### v3

# Uncertainty Quantification with the Empirical Neural Tangent Kernel

Abstract
While neural networks have demonstrated impressive performance across various tasks, accurately quantifying uncertainty in their predictions is essential to ensure their trustworthiness and enable widespread adoption in critical systems. Several Bayesian uncertainty quantification (UQ) methods exist that are either cheap or reliable, but not both. We propose a post-hoc, sampling-based UQ method for overparameterized networks at the end of training. Our approach constructs efficient and meaningful deep ensembles by employing a (stochastic) gradient-descent sampling process on appropriately linearized networks. We demonstrate that our method effectively approximates the posterior of a Gaussian Process using the empirical Neural Tangent Kernel. Through a series of numerical experiments, we show that our method not only outperforms competing approaches in computational efficiency–often reducing costs by multiple factors–but also maintains state-of-the-art performance across a variety of UQ metrics for both regression and classification tasks.
1 Introduction
Neural networks (NN) achieve impressive performance on a wide array of tasks, in areas such as speech recognition (Nassif et al., 2019; Abdel-Hamid et al., 2014; Deng et al., 2013), image classification (LeCun et al., 1998; Krizhevsky et al., 2012; He et al., 2016), computer vision (Redmon et al., 2016; Redmon & Farhadi, 2017), and language processing (Vaswani et al., 2017; Ray, 2023; Devlin et al., 2019), often significantly exceeding human performance. While the promising predictive and generative performance of modern NNs is evident, accurately quantifying uncertainty in their predictions remains an important and active research frontier (Abdar et al., 2021). Models are often over-confident in predictions on out-of-distribution (OoD) inputs (Guo et al., 2017), and sensitive to distribution-shift (Ford et al., 2019). By quantifying a model’s uncertainty, we can determine when it fails to provide well-calibrated predictions, indicating the need for additional training (possibly on more diverse data) or even human intervention. This is vital for deploying NNs in critical applications like diagnostic medicine and autonomous machine control (Nemani et al., 2023b).
An array of uncertainty quantification methods for NNs exist, each with benefits and drawbacks. Frequentist statistical methods, such as conformal prediction (Vovk et al., 2005; Papadopoulos et al., 2002; Lei & Wasserman, 2014), create prediction intervals through parameter estimation and probability distributions on the data. While currently considered state-of-the-art, the main drawback is that conformal prediction is data-hungry, requiring a large hold-out set. Several works Altieri et al. (2024); Dadalto et al. (2023); Granese et al. (2021) employ the feature-space representations of the network to quantify the uncertainty of test predictions, through kernel densities (Kotelevskii et al., 2022), Gaussian Discriminant analysis (Mukhoti et al., 2023), non-constant mapping functions (Tagasovska & Lopez-Paz, 2019), etc. The interpretation of uncertainty in these works is related to risk of misclassification; these methods generally perform very well at detecting OoD points. While both conformal predictions and feature-space works are important, we limit our scope to Bayesian methods in this paper. Though Bayesian methods can suffer from prior misspecification (Masegosa, 2020) and computational burdens, the predictive and posterior distributions of a model are very natural frameworks for quantifying the uncertainty and spread of possible values of the model.
Figure 1: Comparison of various Bayesian UQ methods (see Section˜2) on a 1-layer MLP, trained on the data (red) lying on $y=x^{3}$ (black), with Gaussian noise added. The methods’ mean predictors (blue) $\pm 3\sigma$ (green) are shown, where $\sigma^{2}$ is the variance estimated via each method. We see that NUQLS performs well on this task.
Unfortunately, largely due to the curse of dimensionality, existing Bayesian methods are very expensive to compute, and provide poor approximations to the predictive distribution (Folgoc et al., 2021). This results in a suite of methods that require excessive approximations to scale to large problems (Daxberger et al., 2021), often at the cost of theoretical underpinnings, or necessitating modifications to the network itself (He et al., 2020).
Gaussian Processes (GPs) (Rasmussen & Williams, 2005) are important tools in Bayesian machine learning (ML) that can directly capture the epistemic (model) uncertainty of predictions, and arise as large-width limits of NNs (Neal, 1996). However, naïve GP training scales cubically in the number of training datapoints, necessitating approximations for modern applications. Neural Tangent Kernels (NTKs) (Jacot et al., 2018) describe the functional evolution of a NN under gradient flow, and naturally arise in the analysis of model quality (Hodgkinson et al., 2023). Due to their deep connection to NNs, these covariance functions are enticing as potential tools for UQ of their corresponding NNs.
Motivated by this, we present a UQ method for a trained, over-parameterized NN model, wherein we approximate the predictive distribution through an ensemble of linearized models, trained using (stochastic) gradient-descent. For certain loss functions, this ensemble samples from the posterior of a GP with an empirical NTK.
Contributions. Our contributions are as follows:
  1. 1.
    In Section˜3.1, we present a Monte-Carlo sampling based UQ method to approximate the predictive distribution of NNs, called Neural Uncertainty Quantification by Linearized Sampling (NUQLS). Our method is lightweight, post-hoc, numerically stable, and embarrassingly parallel.
  2. 2.
    Under certain assumptions, Section˜3.2 details the convergence of NUQLS to the predictive distribution of a GP with an empirical NTK kernel, providing a novel perspective on the connection between NNs, GPs, and the NTK.
  3. 3.
    On various ML problems in Section˜4, we show that NUQLS performs as well as or better than leading UQ methods, is less computationally expensive than deep ensemble, and scales to large image classification tasks.
  4. 4.
    In Section˜4.4 we introduce a novel metric for evaluating the quality of UQ methods for classification tasks. This metric more directly measures the quality of UQ methods than existing UQ metrics.
Remark 1.1 (Necessity of Contribution 4.).
Evaluating the quality of UQ methods is non-trivial. There exists in the literature a lack of a suitable framework for evaluating the quality of UQ methods in classification settings. Common metrics such as Negative Log-Likelihood (NLL), Expected Calibration Error (ECE) (Naeini et al., 2015) and Area Under the Curve of the Receiver Operating Characteristic (AUCROC) are actually metrics for prediction quality, or are based on flawed surrogates for uncertainty. Further, it was shown in Abe et al. (2022) that the goal of Bayesian methods should not be to improve predictive ability; performance gains will be attained more economically by choosing a larger model class. Instead, Bayesian methods should seek to accurately quantify the uncertainty of a model, by computing the predictive variance. As uncertainty is not a measurement that can easily be shown to be well-calibrated, one requires a more qualitative approach to evaluating the performance of a UQ model. This is the motivation for the graphical technique we introduce in Section˜4.4 for comparing the quality of UQ estimates for multi-class classification. For an in-depth discussion of these points, please see Appendix˜C.
Notation.
[…]
The process of training amounts to finding a solution, $\widehat{\bm{\theta}}$, to the optimization problem $\min_{\bm{\theta}}\sum_{i=1}^{n}\ell({\bf f}({\bf x}_{i},\bm{\theta}),{\bf y}_{i})+\mathcal{R}(\bm{\theta})$, where $\mathcal{R}(\bm{\theta})$ is a regulariser. For a kernel function ${{\mathbf{K}}}:^{d}\times^{d}\to^{c\times c}$, we define ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}\in^{nc\times nc}$ where the $(i,j)^{\text{th}}$ (block) element is ${{\mathbf{K}}}({\bf x}_{i},{\bf x}_{j})\in^{c\times c}$. Additionally, we define ${{\mathbf{K}}}_{\mathcal{X},{\bf x}}\triangleq\begin{bmatrix}{{\mathbf{K}}}({\bf x}_{1},{\bf x})&\ldots&{{\mathbf{K}}}({\bf x}_{n},{\bf x})\end{bmatrix}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\in^{nc\times c}$ with ${{\mathbf{K}}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}_{\mathcal{X},{\bf x}}={{\mathbf{K}}}_{{\bf x},\mathcal{X}}$. For a matrix $\mathbf{J}$}, its Moore-Penrose pseudo-inverse is denoted by $\mathbf{J}$$^{\dagger}$.
2 Background
Bayesian Framework.
Parametric Bayesian methods admit access to a distribution over predictions ${\bf f}(\bm{\theta},{\bf x}^{\star})$ for unseen test points $\bf x$$^{\star}$, through the posterior distribution $p(\bm{\theta}|\mathcal{D})\propto p(\bm{\theta})p(\mathcal{D}|\bm{\theta})$, and the predictive distribution $p({\bf y}^{\star}|{\bf x}^{\star},\mathcal{D})=\int p({\bf y}^{\star}|{\bf f}(\bm{\theta},{\bf x}^{\star}))p(\bm{\theta}|\mathcal{D})d\bm{\theta}$, where $p(\bm{\theta})$ is the prior, and $p(\mathcal{D}|\bm{\theta})$ is the likelihood function evaluated on the training data. Both the posterior and the predictive distributions are computationally intractable in all but the simplest cases. In our setting, their calculation involves very high-dimensional integrals, which we can approximate through a Monte Carlo (MC) approximation $p({\bf y}^{\star}|{\bf x}^{\star},\mathcal{D})\approx 1/S\sum_{s}p({\bf y}^{\star}|{\bf f}(\bm{\theta}_{s},{\bf x}^{\star}))$ for $\bm{\theta}_{s}\sim q(\bm{\theta})$, where $q(\bm{\theta})$ is an approximation to the posterior distribution. The effectiveness of classical Markov Chain Monte Carlo (MCMC) methods in posterior sampling diminishes in this setting due to the curse of dimensionality, limiting the tractable techniques available with theoretical guarantees. This limitation necessitates coarser approximations for estimating the posterior $q(\bm{\theta})$, leading to the emergence of the following Bayesian methods for posterior approximation.
Proposed in (Gal & Ghahramani, 2016), Monte Carlo Dropout (MC-Dropout) takes a trained NN, and uses the dropout regularization technique at test time to sample $S$ sub-networks, $\left\{{\bf f}(\bm{\theta}_{s},{\bf x})\right\}_{s=1:S}$ as an MC approximation of the predictive distribution. MC-Dropout is an inexpensive method, yet it is unlikely to converge to the true posterior, and is erroneously multi-modal (Folgoc et al., 2021).
In Variational Inference (VI) (Hinton & Van Camp, 1993; Graves, 2011), a tractable family of approximating distributions for $p(\bm{\theta}|\mathcal{D})$ is chosen, denoted by $q_{\bm{\psi}}(\bm{\theta})$, and parameterized by $\bm{\psi}$. The optimal distribution in this family is obtained by finding $\bm{\psi}$ that minimizes the Kullback-Leibler divergence between $q_{\bm{\psi}}(\bm{\theta})$ and $p(\bm{\theta}|\mathcal{D})$. To be computationally viable, mean field and low-rank covariance structures are often required for $q_{\bm{\psi}}(\bm{\theta})$.
The Laplace Approximation (LA) (MacKay, 1992; Ritter et al., 2018) is a tool from classical statistics which approximates the posterior distribution by a Gaussian centered at the maximum a posteriori solution (MAP) with normalized inverse Fisher information covariance. This is justified by the Bernstein-von Mises Theorem (Van der Vaart, 2000, pp. 140–146), which guarantees that the posterior converges to this distribution in the large-data limit, for well-specified regular models. However, NNs are often over-parameterized, and the regime where $n\to\infty$ with fixed $p$ is no longer valid or a reasonable reflection of modern deep learning models (De Bortoli & Desolneux, 2022).
These limitations are acknowledged but seldom discussed by the Bayesian Deep Learning (BDL) community, which tends to view this as an additional layer of approximation rather than a modelling error, leading to the development of synonymous LA-inspired methods. However, we will show that such methods typically perform poorly compared to deep ensembles, which are often excluded from comparisons.
We can evaluate the posterior and predictive distribution in the LA using the linearization of ${\bf f}({\bf x},\bm{\theta})$ around the MAP solution. This approach is known as the Linearized Laplace Approximation (LLA) and typically delivers better performance than LA (Immer et al., 2021). LLA generally requires reduction to a subset of parameters or approximations of the covariance structure (Martens & Grosse, 2015) to scale, at the cost of performance. Recent work (Antorán et al., 2022; Ortega et al., 2023) has enabled LLA to become more scalable with better performance for larger models and datasets.
Deep Ensembles (DE) (Lakshminarayanan et al., 2017) are comprised of $S$ networks that are independently trained on the same training data, with different initializations, leading to a collection of parameters $\{\bm{\theta}_{s};s=1\ldots,S\}$. At test time, our predictive distribution becomes $p({\bf y}|{\bf x},\mathcal{D})\approx 1/S\sum_{s=1}^{S}p({\bf y}|{\bf f}(\bm{\theta}_{s},{\bf x}))$. Despite their simple construction, DEs are able to obtain samples from different modes of the posterior, and are often considered state-of-the-art for BDL (Hoffmann & Elster, 2021). However, due to the often large cost of training performant neural networks, deep ensembles of reasonable size can be undesirably expensive to obtain.
Stochastic Weight Averaging Gaussian (SWAG) (Maddox et al., 2019) takes a trained network and undergoes further epochs of SGD training to generate a collection of parameter samples. A Gaussian distribution with sample mean and a low-rank approximation of the sample covariance is then used to approximate a posterior mode.
Gaussian Processes.
A GP is a stochastic process that is defined by a mean and a kernel function. A GP models the output of a random function ${\bf f}:^{d}\to^{c}$, at a finite collection of points $\bf x$}, as being jointly Gaussian distributed. Conditioning on training data $\mathcal{D}$, it generates a posterior predictive distribution $p({\bf f}({\bf x}_{\star})|\mathcal{D})$ at a test point $\bf x$$_{\star}$. For example, in regression settings where $c=1$, with the mean and kernel functions $\mu:^{d}\to$ and $\kappa:^{d}\times^{d}\to$, as well as the observations $y\sim\mathcal{N}(f({\bf x}),\sigma^{2})$, there is a closed form expression for the predictive distribution, $p(f({\bf x}_{\star})|\mathcal{D})\sim\mathcal{N}\big(\bm{\mu}({\bf x}_{\star}),\bm{\sigma}({\bf x}_{\star})\big)$, where
$\bm{\mu}({\bf x}_{\star})$ $={{\bf k}}_{{\bf x}_{\star},\mathcal{X}}\big[{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}+\sigma^{2}\mathbf{I}\big]^{-1}\big({\bf y}-\bm{\mu}(\mathcal{X})\big)+\bm{\mu}({\bf x}_{\star})$
$\bm{\sigma}({\bf x}_{\star})$ $=\kappa({\bf x}_{\star},{\bf x}_{\star})-{{\bf k}}_{{\bf x}_{\star},\mathcal{X}}\big[{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}+\sigma^{2}\mathbf{I}\big]^{-1}{{\bf k}}_{\mathcal{X},{\bf x}_{\star}}$
for ${\bf y}\triangleq[y_{1},\dots,y_{n}]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}$ and $\bm{\mu}(\mathcal{X})\triangleq[\mu({\bf x}_{1}),\dots,\mu({\bf x}_{n})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}$. GPs can yield impressive predictive results when a suitable kernel is chosen (Rasmussen, 1997). However, forming the kernel and solving linear systems makes GP computations intractable for large datasets. Approximations such as sparse variational inference (Titsias, 2009), Nyström methods (Martinsson & Tropp, 2020), and other subspace approximations (Gardner et al., 2018) can alleviate the computational burden; however, these approximations often result in a significant decline in predictive performance.
Neural Tangent Kernel.
Under continuous time gradient flow, it can be shown that a NN output ${\bf f}(\cdot,\bm{\theta}):^{d}\to^{c}$ undergoes kernel gradient descent, namely $\partial_{t}{\bf f}({\bf x},\bm{\theta}_{t})=-\sum_{i=1}^{n}{{\mathbf{K}}}_{\bm{\theta}_{t}}({\bf x},{\bf x}_{i})\nabla_{{\bf f}}\ell({\bf f}({\bf x}_{i},\bm{\theta}_{t}),{\bf y}_{i})$, where
${{\mathbf{K}}}_{\bm{\theta}}({\bf x},{\bf y})\triangleq\left\langle\frac{\partial{\bf f}}{\partial\bm{\theta}}({\bf x},\bm{\theta}),\frac{\partial{\bf f}}{\partial\bm{\theta}}({\bf y},\bm{\theta})\right\rangle\in^{c\times c},$ (1)
is the empirical NTK (Jacot et al., 2018). As the width of a network increases, the empirical NTK converges (in probability) to a deterministic limit, sometimes referred to as the analytic NTK, that is independent of the network’s parameters. Lee et al. (2019) showed that in this limit, the network acts according to its NTK linearization during gradient descent (GD) training. This parameter independence results in a loss of feature learning in the limiting regime (Yang & Hu, 2021). However, for finite-width NNs, Fort et al. (2020) empirically showed that the empirical NTK becomes “data-dependent” during training. Since we focus exclusively on the finite-width regime, we refer to the empirical NTK simply as the NTK.
Related Works
Our method NUQLS shares notable similarities with, and exhibits distinct differences from, several prior works, such as Sampling-LLA, Bayesian Deep Ensembles and local ensembles. Due to the breadth and depth of this discussion, we relegate the related works discussion to Appendix˜B.
3 NUQLS
Algorithm 1 NUQLS
Input: number of realizations $S$, weights $\widehat{\bm{\theta}}$.
for $s=1$ to $S$ do
  $\bm{\theta}_{0,s}\leftarrow\widehat{\bm{\theta}}+{\bf z}_{0}$, where ${\bf z}_{0}\sim\mathcal{N}({\bf 0},\gamma^{2}\mathbf{I})$
  $\bm{\theta}^{\star}_{s}\leftarrow$ Run (stochastic) GD from $\bm{\theta}_{0,s}$ to
    (approximately) solve ˜3 and obtain $\bm{\theta}^{\star}_{s}$
end for
return $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},.)\}_{s=1}^{S}$
We now present Neural Uncertainty Quantification by Linearized Sampling (NUQLS), our post-hoc sampling method for quantifying the uncertainty of a trained NN. We begin by presenting the motivation and a high-level overview of our method. Subsequently, we provide theoretical justification, demonstrating that, under specific conditions, the NUQLS samples represent draws from the approximate posterior of the neural network, which is equivalent to a GP defined by the NTK.
3.1 Motivation and High-level Overview
NNs are often over-parameterized, resulting in non-uniqueness of interpolating solutions, with sub-manifolds of parameter space able to perfectly predict the training data (Hodgkinson et al., 2023). To generate a distribution over predictions, we adopt a Bayesian framework, where the uncertainty in a neural network’s prediction can be interpreted as the spread of possible values the network might produce for a new test point, conditioned on the training data. To quantify this uncertainty, we can evaluate the test point on other “nearby” models with high posterior probability and analyze their range of predictions. To identify such models, we propose using the linearized approximation of the original network around its trained parameters as a simpler yet expressive surrogate. This approach can retain, to a great degree, the rich feature representation of the original network while enabling tractable exploration of the posterior distribution. In the same spirit as DE, in the overparameterized setting, we can fit this linear model to the original training data, using (stochastic) gradient descent with different initializations, resulting in an ensemble of linear predictors. Not only does this ensemble explain the training data well, but it also provides a practical way to estimate predictive uncertainty.
More precisely, let $\widehat{\bm{\theta}}$ be a set of parameters obtained after training the original NN. Linearizing $\bf f$} around $\widehat{\bm{\theta}}$ gives
[…]
where ${\mathbf{J}}(.,{\bf x})=[\partial{\bf f}(.,{\bf x})/\partial\bm{\theta}]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\in^{c\times p}$ is the Jacobian of $\bf f$}. Using the linear approximation ˜2, we consider
$\min_{\bm{\theta}}\;$ $\sum_{i=1}^{n}\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i}).$ (3)
In overparameterized settings, ˜3 may have infinitely many solutions. To identify these solutions and create our ensemble of linear predictors, we employ (stochastic) gradient descent, initialized at zero-mean isotropic Gaussian perturbations of the trained parameter, $\widehat{\bm{\theta}}$. The pseudo-code for this algorithm is provided in Algorithm˜1. For a given test point $\bf x$$^{\star}$}, the mean prediction and uncertainty can be computed using $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x}^{\star})\}_{s=1}^{S}$.
Note that while the training cost for a linearised network is only slightly higher per epoch compared to standard NN training, each network in the NUQLS ensemble is initialized in a neighborhood of a local minimum of the original NN. As a result, NUQLS often requires significantly fewer epochs to converge, leading to an order-of-magnitude computational speedup relative to DE (see Tables˜1, 2 and 9 for wall-clock time comparisons, and Appendix˜E for a more in-depth analysis of the computation costs.).
3.2 Theoretical Analysis
We now establish the key property of Algorithm˜1: under mild conditions, NUQLS generates samples from the approximate posterior of the neural network, which in many cases corresponds to a Gaussian process defined by the NTK. Proofs are provided in Appendix˜A.
Suppose $\bm{\theta}^{\ddagger}$ is any solution to ˜3. Using $\bm{\theta}^{\ddagger}$, one can construct a family of solutions to ˜3 as
$\bm{\theta}^{\star}_{{\bf z}}=\bm{\theta}^{\ddagger}+\left(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{\dagger}{\mathbf{J}}_{\mathcal{X}}\right){\bf z},\quad\forall{\bf z}\in^{p},$ (4)
[…]
$\min_{\bm{\theta}}\;$ $\sum_{i=1}^{n}\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i}),\;\;\text{s.t.}\;\;\bm{\theta}\in{\textnormal{Range}}\left({\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right).$ (5)
Lemma 3.1.
Suppose the loss, $\ell(\,\cdot\,,{\bf y})$, is either:
  • strongly convex in its first argument, or
  • strictly convex in its first argument, and a solution to ˜3 exists.
The problem ˜5 admits a unique solution.
As it turns out, any solution of the form ˜4 can be efficiently obtained using (stochastic) gradient descent.
Theorem 3.2.
Consider the optimization problem ˜3 and assume ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank.
  • (Gradient Descent) Suppose $\ell({\bf f},{\bf y})$ is strictly convex with locally Lipschitz continuous gradient, both with respect to $\bf f$}, and the problem ˜3 admits a solution. Gradient descent, initialized at $\bf z$} and with appropriate learning rate, converges to ˜4.
  • (Stochastic Gradient Descent) Suppose $\ell({\bf f},{\bf y})$ is strongly convex with Lipschitz continuous gradient, both with respect to $\bf f$}, and any solution to the problem ˜3 is interpolating. Stochastic gradient descent, initialized at $\bf z$} and with small enough learning rate, converges to ˜4 with probability one.
We note that the local smoothness requirement in the first part of Theorem˜3.2 is a relatively mild assumption; for example, it holds if we simply assume that $\ell$ is twice continuously differentiable. Also, the full-row rank assumption on ${\mathbf{J}}_{\mathcal{X}}$ in the second part of Theorem˜3.2 is reasonable for highly over-parameterized networks; e.g., see Liu et al. (2022).
Now, suppose ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank and the assumption of Lemma˜3.1 holds, ensuring the existence of the unique solution $\bm{\theta}^{\ddagger}$ to ˜5. Noting ${\textnormal{Range}}\left({\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)={\textnormal{Range}}({\mathbf{J}}_{\mathcal{X}}^{\dagger})$, we can write ˜4 as
$\bm{\theta}^{\star}_{{\bf z}}=$ ${\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\bf w}+\left(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}\right){\bf z},\quad\forall{\bf z}\in^{p}$
where $\bf w$} is a vector for which $\bm{\theta}^{\ddagger}={\mathbf{J}}_{\mathcal{X}}^{\dagger}{\bf w}$, and ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}\triangleq{\mathbf{J}}_{\mathcal{X}}{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}_{\mathcal{X}}={{\mathbf{K}}}_{\widehat{\bm{\theta}}}(\mathcal{X},\mathcal{X})\in^{nc\times nc}$ is the Gram matrix of the empirical NTK ˜1 on the training data $\mathcal{X}$. Setting ${\bf z}=\widehat{\bm{\theta}}-{\bf z}_{0}$ for some ${\bf z}_{0}$, we get
$\bm{\theta}^{\star}_{{\bf z}}$ $={\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\bf w}+\left(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}\right)(\widehat{\bm{\theta}}-{\bf z}_{0}),$
$\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})$ $={\bf f}(\widehat{\bm{\theta}},{\bf x})+{{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}\left({\bf w}-{\mathbf{J}}_{\mathcal{X}}\widehat{\bm{\theta}}\right)+\left({{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}-{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x})\right){\bf z}_{0},$ (6)
where ${{\mathbf{K}}}_{{\bf x},\mathcal{X}}\triangleq{{\mathbf{K}}}_{\widehat{\bm{\theta}}}({\bf x},\mathcal{X})\in^{c\times nc}$. Taking ${\bf z}_{0}$ to be a random variable, we form an ensemble of predictors $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\}_{{\bf z}}$, where each $\bm{\theta}^{\star}_{{\bf z}}$ is formed from the projection of a random $\bf z$} onto $\text{Null}({\mathbf{J}}_{\mathcal{X}})$. We require $\bf z$}’s distribution to be symmetric, isotropic, and centered at $\widehat{\bm{\theta}}$, as we do not know a priori which directions contain more information. We take the maximum entropy distribution for a given mean and variance, which is Gaussian [^1] Heavier-tailed distributions matching the above criteria, e.g. logistic distributions, may improve results.. Hence, we let ${\bf z}_{0}\sim\mathcal{N}({\bf 0},\gamma^{2}\mathbf{I})$, for some hyper-parameter $\gamma\in$. The expectation and variance of ˜6, and the distribution of the predictor ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})$, are then
${\mathbb{E}}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\big)=\mathbf{\bm{\mu}}(\widehat{\bm{\theta}},{\bf x})$ $={{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}(\bm{\theta}^{\ddagger}-\widehat{\bm{\theta}})+{\bf f}(\widehat{\bm{\theta}},{\bf x}),$ (7)
$\text{Var}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\big)=\mathbf{\bm{\sigma}}^{2}(\widehat{\bm{\theta}},{\bf x})$ $=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}.$ (8)
${\bf f}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\overset{\text{approx}}{\sim}$ $\mathcal{N}\big(\bm{\mu}(\widehat{\bm{\theta}},{\bf x}),\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})\big).$ (9)
Remark 3.3 (Connections to GP: Regression).
[…]
$\mu(\widehat{\bm{\theta}},{\bf x})$ $={{\bf k}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}({\bf y}-{\bf f}(\widehat{\bm{\theta}},\mathcal{X}))+f(\widehat{\bm{\theta}},{\bf x}),$
$\sigma^{2}(\widehat{\bm{\theta}},{\bf x})$ $=\big(\kappa({\bf x},{\bf x})-{{\bf k}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\bf k}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}.$
By the full-rank assumption on the Jacobian, this amounts to the conditional distribution of the following normal distribution, conditioned on interpolation ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf y}$,
[…]
Therefore, $f(\bm{\theta}^{\star}_{{\bf z}},{\bf x})$ follows a GP with an NTK kernel. Conditioning on ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf y}$ is reasonable, since by construction ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})\approx\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})$, and under the full-rank assumption of ${\mathbf{J}}_{\mathcal{X}}$, we have $\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf y}$.
Remark 3.4 (Connections to GP: General Loss).
Beyond quadratic loss, for a general loss function satisfying the assumptions of Lemma˜3.1, a clear GP posterior interpretation like that in ˜10 may not exist. Nevertheless, we can still derive related insights. If $\widehat{\bm{\theta}}$ is an interpolating solution from initial training, which is common for modern NNs, then as long as ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank, solving ˜3 is equivalent to finding $\bm{\theta}\in^{p}$ such that $\widetilde{{\bf f}}(\bm{\theta},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., find $\bm{\theta}\in^{p}$ such that ${\mathbf{J}}_{\mathcal{X}}(\bm{\theta}-\widehat{\bm{\theta}})={\bf 0}$. So we obtain (9) with $\bm{\mu}(\widehat{\bm{\theta}},{\bf x})={\bf f}(\widehat{\bm{\theta}},{\bf x})$ and $\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}$ as the conditional distribution of ˜10, conditioned on the event ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., interpolation with $\ell({\bf f}(\bm{\theta}^{\star}_{{\bf z}},{\bf x}_{i}),{\bf y}_{i})=0$ for $i=1,\ldots,n$.
The Punchline.
Drawing samples from the posterior ˜9 by explicitly calculating ˜7 and 8 can be intractable in large-scale settings. Moreover, it can be numerically unstable due to the highly ill-conditioned nature of the NTK matrix [^2] Recall that the condition number of ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}$ is the square of that of ${\mathbf{J}}_{\mathcal{X}}$.. However, by combining the above derivations with Theorem˜3.2, we arrive at the key property of Algorithm˜1: it enables efficient sampling from the posterior ˜9.
Corollary 3.5 (Key Property of NUQLS).
With the assumptions of Theorem˜3.2, the samples generated by Algorithm˜1 represent draws from the predictive distribution in ˜9.
Hence, we approximate ˜7 and 8 by computing the sample mean and covariance of $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x})\}_{s=1}^{S}$ obtained from Algorithm˜1. By the law of large numbers, the quality of these approximations improves as $S\to\infty$.
Remark 3.6.
Loss functions that do not satisfy the assumption of Lemma˜3.1, such as the cross-entropy loss, may fail to yield a unique representation of ˜4, so the above posterior analysis does not apply. However, our experiments demonstrate that Algorithm˜1 can still generate samples that effectively capture the posterior variance (see Section˜4.4) and posterior mean (see Section˜G.1). In cases where ˜3 lacks a solution, Algorithm˜1 can still be executed by terminating the iterations of (stochastic) GD early. Investigating the distribution of the resulting ensemble and its connection to an explicit posterior remains a potential direction for future research.
4 Experiments
We now empirically demonstrate the result of Theorem˜3.2, as well as compare the performance of our method with alternatives on various regression and classification tasks. Implementation details are given in Appendix˜I. The PyTorch implementation of our experiments is available here. We have also released our method as a package. For additional experimental results, please see Appendix˜H.
4.1 Empirical Convergence
[…]
We demonstrate empirically the convergence of the predictive distribution of NUQLS to that of an NTK-GP. We take a series of MLPs with NTK scaling (Jacot et al., 2018), and train these NNs on normalized Gaussian data (a regression task). Under this setting, the condition number of the resultant NTK is reasonable ($\approx 10^{3}-10^{8}$, compared to $\approx 10^{17}$ on some UCI regression datasets). This allows us to explicitly compute the predictive variance for the NTK-GP for some normalized Gaussian test set. We then also compute the predictive variance using NUQLS, and take the $l_{2}$ norm of the difference between the two variance sets, which we term Squared Error of the Variance (SEV). We plot this against number of epochs of training for NUQLS, and the number of NUQLS realizations $s$. We also plot the average training loss of the linearised networks in the NUQLS ensemble. The values are averaged over $10$ random realizations. The results are displayed in Figure˜2. Note that the SEV cannot be zero, due to the condition number of the NTK; this affects the solution of the linear system $K_{XX}\mathbf{y}=K_{Xx}$ in the computation of the variance for the NTK-GP. However, we see clear convergence of our method to the distribution of an NTK-GP as the ensemble members of NUQLS approach their minima, and as the number of ensemble members increases.
4.2 Toy Regression
We compare the performance of our method on a toy regression problem, taken from (Hernández-Lobato & Adams, 2015) and extended in (Park & Blei, 2024). In Figure˜1, we take $20$ uniformly sampled points in the domain $x\in[-4,-2]\cup[2,4]$, and let $y=x^{3}+\epsilon,~\epsilon\sim\mathcal{N}(0,3^{2})$. A small MLP was trained on these data and used for prediction. We apply VI, SWAG, LA, LLA, DE and NUQLS to the network to find a predictive mean and uncertainty. Close to the training data, i.e. in $4,-2]\cup[2,4]$ we expect low uncertainty; outside of this region, the uncertainty should grow with distance from the training points. VI underestimates, while SWAG and LA overestimate the uncertainty. DE grows more uncertain with distance from the training points, however both NUQLS and the LLA contain the underlying target curve within their confidence intervals. Note that deep ensembles output a heteroskedastic variance term, and were trained on a Gaussian likelihood; in comparison, the variances for LLA and NUQLS were computed post-hoc.
4.3 UCI Regression
Table 1: Comparing performance of NUQLS, DE, LLA and SWAG on UCI regression tasks. NUQLS performs as well as or better than all other methods, while showing a speed up over other methods; this speed up increases with the size of the datasets. LLA-K denotes LLA with a KFAC covariance structure. Reported time for NUQLS, LLA and SWAG includes the training time for the original NN, with the run-time of the post-hoc method given in brackets.
Dataset Method RMSE $\downarrow$ NLL $\downarrow$ ECE $\downarrow$ Time(s)
Energy NUQLS $\mathbf{0.047\pm}$$\mathbf{0.006}$ $\mathbf{-2.400\pm}$$\mathbf{0.209}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $\mathbf{8.374}~(\mathbf{0.151})$
DE $218\pm0.0$320.032 $\mathbf{-1.651\pm}$$\mathbf{0.783}$ $\mathbf{0.004\pm}$$\mathbf{0.002}$ $102.244$
LLA $\mathbf{0.048\pm}$$\mathbf{0.006}$ $\mathbf{-2.475\pm}$$\mathbf{0.128}$ $\mathbf{0.004\pm}$$\mathbf{0.004}$ 8.491(0.269)8.491~(0.269)
SWAG $\mathbf{0.058\pm}$$\mathbf{0.015}$ $1.950\pm$$0.158$ $080\pm0.0$110.011 45.306(37.084)45.306~(37.084)
Kin8nm NUQLS $\mathbf{0.252\pm}$$\mathbf{0.005}$ $796\pm0.0$250.025 $\mathbf{0.000\pm}$$\mathbf{0.000}$ $\mathbf{26.570}~(\mathbf{0.264})$
DE $\mathbf{0.252\pm}$$\mathbf{0.006}$ $\mathbf{-0.914\pm}$$\mathbf{0.028}$ $002\pm0.0$010.001 $73.967$
LLA $\mathbf{0.260\pm}$$\mathbf{0.010}$ $783\pm0.0$540.054 $\mathbf{0.001\pm}$$\mathbf{0.001}$ 38.272(11.966)38.272~(11.966)
SWAG $0.457\pm$$0.149$ $0.006\pm$$0.295$ $054\pm0.0$120.012 176.569(150.263)176.569~(150.263)
Protein NUQLS $\mathbf{0.623\pm}$$\mathbf{0.005}$ $\mathbf{0.209\pm}$$\mathbf{0.047}$ $\mathbf{0.002\pm}$$\mathbf{0.000}$ $\mathbf{81.264}~(\mathbf{1.356})$
DE $741\pm0.0$520.052 $\mathbf{0.203\pm}$$\mathbf{0.203}$ $\mathbf{0.011\pm}$$\mathbf{0.020}$ $1014.827$
LLA-K $640\pm0.0$070.007 $458\pm0.0$710.071 $\mathbf{0.002\pm}$$\mathbf{0.000}$ 89.414(9.506)89.414~(9.506)
SWAG $730\pm0.0$440.044 $\mathbf{0.187\pm}$$\mathbf{0.080}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ 548.88(468.972)548.88~(468.972)
In Tables˜1 and 9, we compare NUQLS with DE, LLA and SWAG on a series of UCI regression problems. Mean squared error (MSE) and expected calibration error (ECE) respectively evaluate the predictive and UQ performance, with Gaussian negative log likelihood (NLL) evaluating both. See (Nemani et al., 2023a, $\S 4.11$) for an explanation of ECE. We see that NUQLS consistently has the (equal) best ECE (except for the Song dataset, where it falls short of the LLA ECE by $\mathit{0.1\%}$). It has comparable or better NLL than other methods on all datasets, and often gives an improvement on RMSE. Finally, it is the quickest method, often by a very significant margin, and it does not fail on any datasets, like the other methods do. Note that for the two largest datasets, Protein and Song, we required approximations on the covariance structure of LLA (see (Daxberger et al., 2021)). A detailed explanation of the hyper-parameter tuning method for NUQLS is given in Section˜F.1.
4.4 Image Classification - Uncertainty
Figure 3: Violin plot of VMSP, for correctly predicted ID test points, incorrectly predicted ID test points, and OoD test points. Median is shown, with violin width depicting density. Low variance is expected for ID correct points, and large variance for ID incorrect and OoD points. Title of plots gives model and dataset used for training.
We now compare the UQ performance of NUQLS, DE, SWAG, LLA* (LLA with a last-layer and KFAC approximation), and MC-Dropout (MC), on larger image classification tasks. We take variance of the maximum predicted softmax probability (VMSP), for a given test point, as the correct quantifier of uncertainty in this setting (see Appendix˜C for justification).
Figure˜3 presents a violin plot of the VMSP for three test-groups: correctly predicted in-distribution (FashionMNIST, CIFAR-10, CIFAR-100) test points, incorrectly predicted in-distribution test points, and out-of distribution (MNIST, CIFAR-100, CIFAR-10) test points. We would expect that there should be, on average, much smaller uncertainty for ID test points that have been correctly predicted, and larger uncertainty for incorrectly predicted ID and OoD test points. We compare against a completely randomized baseline method (BASE), where we sample $10$ standard normal realizations of logits, passed through a softmax. In Table˜10 in Section˜H.8, we display the corresponding median and sample skew values for each method in each test group, to quantify the distribution of VMSP for each test set. Ideally, a method should far outperform the baseline, so we can use the median and sample skew difference between a method and the baseline as a way to compare different methods. We see that NUQLS outperforms all other methods, including the SOTA method DE. In Section˜H.7 we provide additional experimental evaluation of NUQLS. In Figure˜6 we evaluate NUQLS on a ResNet50 trained on both SVHN and ImageNet, displaying the scalability of our method, as well as providing comparison with other competing methods, including Bayesian Deep Ensembles (BDE), Spectral-Normalized Neural Gaussian Process (SNGP), BatchEnsemble (BE), and Stochastic Gradient Langevin Dynamics (SGLD). Against these competing methods, NUQLS performs the strongest.
5 Conclusion
We have presented NUQLS, a Bayesian, post-hoc UQ method that approximates the predictive distribution of an over-parameterized NN through GD/SGD training of linear networks, allowing scalability without sacrificing performance. Under assumptions on the loss function, this predictive distribution reduces to a GP using the NTK. We find that our method is competitive with, and often far outperforms, existing UQ methods on regression and classification tasks, whilst providing a novel connection between NNs, GPs and the NTK.
Limitation.
A theoretical limitation of this work is that its connection to the NTK-GP does not extend to loss functions that violate the assumptions of Theorem˜3.2, such as the cross-entropy loss. The strong empirical performance of NUQLS on classification tasks motivates future research to extend Theorem˜3.2 to broader classes of loss functions and alternative optimization algorithms. A further limitation of the method is the dependence of NUQLS on the linearization approximation. As can be seen in Section˜H.1, when the target neural network is poorly trained, and hence the loss-landscape is far from flat around the ’trained’ parameters, the performance of NUQLS suffers. However, when the network is well-trained, as is common in practical settings, the linear approximation holds and hence NUQLS performs well.
Acknowledgments and Disclosure of Funding
Liam Hodgkinson is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144). Fred Roosta was partially supported by the Australian Research Council through an Industrial Transformation Training Centre for Information Resilience (IC200100022).
Appendix A Proofs
Proof of Lemma˜3.1.
First, we note that by the assumption on $\ell$, a solution to ˜5 always exists. Suppose to the contrary that ˜5 has two distinct solutions $\tilde{\bm{\theta}}$ and $\widehat{\bm{\theta}}$ such that $\tilde{\bm{\theta}}\neq\widehat{\bm{\theta}}$. Since $\tilde{\bm{\theta}}\in{\textnormal{Range}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}})$ and $\widehat{\bm{\theta}}\in{\textnormal{Range}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}})$, i.e., $(\tilde{\bm{\theta}}-\widehat{\bm{\theta}})\perp{\textnormal{Null}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})])$, it follows that $^{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\tilde{\bm{\theta}}\neq{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\widehat{\bm{\theta}}$, which in particular implies $\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\tilde{\bm{\theta}}\rangle\neq\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\widehat{\bm{\theta}}\rangle$ for all $i=1,\ldots,n$. Consider $\bar{\bm{\theta}}=(\tilde{\bm{\theta}}+\widehat{\bm{\theta}})/2$. By strict convexity on $\ell$, we have
$\sum_{i=1}^{n}\ell(\widetilde{{\bf f}}(\bar{\bm{\theta}},{\bf x}_{i}),y_{i})$ $=\sum_{i=1}^{n}\ell\Bigg({\bf f}(\widehat{\bm{\theta}},{\bf x}_{i})+\left\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\frac{\tilde{\bm{\theta}}+\widehat{\bm{\theta}}}{2}-\widehat{\bm{\theta}}\right\rangle,y_{i}\Bigg)$
$=\sum_{i=1}^{n}\ell\Bigg(\frac{{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i})+\left\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\tilde{\bm{\theta}}-\widehat{\bm{\theta}}\right\rangle}{2}+\frac{{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i})+\left\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\widehat{\bm{\theta}}-\widehat{\bm{\theta}}\right\rangle}{2},y_{i}\Bigg)$
$<\frac{1}{2}\sum_{i=1}^{n}\ell\left(\widetilde{{\bf f}}(\tilde{\bm{\theta}},{\bf x}_{i}),y_{i}\right)+\frac{1}{2}\sum_{i=1}^{n}\ell\left(\widetilde{{\bf f}}(\widehat{\bm{\theta}},{\bf x}_{i}),y_{i}\right)$
$=\sum_{i=1}^{n}\ell\left(\widetilde{{\bf f}}(\tilde{\bm{\theta}},{\bf x}_{i}),y_{i}\right),$
which is a contradiction. ∎
Proof of Theorem˜3.2.
  • (Gradient Descent) Denoting
[…]
    we can write
    ${\bf z}={\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X}){\bf z}+\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z},$
    where $\widehat{\bm{\theta}}$ represents the parameters around which the linear model $\widetilde{{\bf f}}$ is defined in (2). The first iteration of gradient descent, initialized at $\bm{\theta}^{(0)}={\bf z}$, is given by
    $\bm{\theta}^{(1)}$ $={\bf z}-\alpha\sum_{i=1}^{n}\frac{\partial\widetilde{{\bf f}}}{\partial\bm{\theta}}({\bf z},{\bf x}_{i})\nabla\ell(\widetilde{{\bf f}}({\bf z},{\bf x}_{i}),{\bf y}_{i})$
    $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X}){\bf z}-\alpha\sum_{i=1}^{n}\frac{\partial\widetilde{{\bf f}}}{\partial\bm{\theta}}({\bf z},{\bf x}_{i})\nabla\ell(\widetilde{{\bf f}}({\bf z},{\bf x}_{i}),{\bf y}_{i})$
    $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X}){\bf z}-\alpha\sum_{i=1}^{n}\left[{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\nabla\ell(\widetilde{{\bf f}}({\bf z},{\bf x}_{i}),{\bf y}_{i})$
    $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\bf v}^{(1)},$
    where ${\bf v}^{(1)}\in{\textnormal{Range}}\left(\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)$. The next iteration is similarly given by
    $\bm{\theta}^{(2)}$ $=\bm{\theta}^{(1)}-\alpha\sum_{i=1}^{n}\frac{\partial\widetilde{{\bf f}}}{\partial\bm{\theta}}(\bm{\theta}^{(1)},{\bf x}_{i})\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{(1)},{\bf x}_{i}),{\bf y}_{i})$
    $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\bf v}^{(1)}-\alpha\sum_{i=1}^{n}\left[{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{(1)},{\bf x}_{i}),{\bf y}_{i})$
    $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\bf v}^{(1)}+{\bf v}^{(2)},$
    where again ${\bf v}^{(2)}\in{\textnormal{Range}}\left(\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)$. Generalizing to the $n$th iteration,
    $\bm{\theta}^{(n)}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\sum_{i=1}^{n}{\bf v}^{(i)}.$
    Hence, by the assumption on $\ell$, as long as an adaptive learning rate is chosen appropriately according to Malitsky & Mishchenko (2020), GD must converge to a solution of the form
    $\bm{\theta}^{\star}_{{\bf z}}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\sum_{i=1}^{\infty}{\bf v}^{(i)}$
    $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\bf v},$
    where ${\bf v}\in{\textnormal{Range}}\left(\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)$. In particular, for ${\bf z}={\bf 0}$, by Lemma˜3.1, we must have that ${\bf v}=\bm{\theta}^{\ddagger}$, where $\bm{\theta}^{\ddagger}$ is the solution to ˜5. Therefore,
    $\bm{\theta}^{\star}_{{\bf z}}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\bm{\theta}^{\ddagger}.$
  • (Stochastic Gradient Descent) Using a similar argument as above, it is easy to show that each iteration of the mini-batch SGD is of the form
    $\bm{\theta}^{(n)}\in\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+{\textnormal{Range}}\left(\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right).$
    Hence, it suffices to show that SGD converges almost surely. Defining
    $\mathcal{L}(\bm{\theta})\triangleq\sum_{i=1}^{n}\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i}),\quad\text{and}\quad{\bf g}(\bm{\theta},\mathcal{X})\triangleq\begin{bmatrix}\nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{1}),{\bf y}_{1})\\ \vdots\\ \nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{n}),{\bf y}_{n})\end{bmatrix}\in^{nc},$
    we write
    $\nabla\mathcal{L}(\bm{\theta})=\sum_{i=1}^{n}\left[{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{1})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})=\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{\bf g}(\bm{\theta},\mathcal{X}).$
    Let $\bm{\theta}^{\star}_{{\bf z}}$ be any solution to ˜3. The full row-rank assumption on ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})$ implies that we must have ${\bf g}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf 0}$, i.e., $\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x}_{i}),{\bf y}_{i})={\bf 0}$, $i=1,\ldots,n$. By the $\mu$-strong convexity assumption on $\ell(.,{\bf y})$ with respect to its first argument, it is easy to see that $\mathcal{L}(\bm{\theta})$ satisfies the Polyak-Łojasiewicz inequality (Karimi et al., 2016) with constant $2\mu\lambda$ where
    $\lambda\triangleq\min_{i=1,\ldots,n}\sigma^{2}_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})),$
    and $\sigma_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i}))$ is the smallest non-zero singular value of ${\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})$. Indeed, let $\bm{\theta}^{\star}_{{\bf z}}$ be any solution to ˜3. From $\mu$-strong convexity of $\ell(.,{\bf y})$ with respect to its first argument, for any $\bm{\theta}$, we have
    $\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})-\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x}_{i}),{\bf y}_{i})$ $\leq\frac{1}{2\mu}\left\|\nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})\right\|^{2}$
    $\leq\frac{1}{2\mu\sigma_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i}))}\left\|\left[{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})\right\|^{2},$
    which implies
    $\mathcal{L}(\bm{\theta})-\mathcal{L}(\bm{\theta}^{\star})$ $\leq\frac{1}{2\mu\lambda}\left\|\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{\bf g}(\bm{\theta},\mathcal{X})\right\|^{2}=\frac{1}{2\mu\lambda}\left\|\nabla\mathcal{L}(\bm{\theta})\right\|^{2}.$
    By the smoothness assumption on each $\ell(\widetilde{{\bf f}}(.,{\bf x}_{i}),{\bf y}_{i})$ as well as the interpolating property of $\bm{\theta}^{\star}_{{\bf z}}$, Bassily et al. (2018, Theorem 1) implies that the mini-batch SGD with small enough step size $\eta$ has an exponential convergence rate as
    ${\mathbb{E}}\mathcal{L}(\bm{\theta}^{(k)})$ $\leq(1-\rho)^{k}\mathcal{L}(\bm{\theta}^{(0)}),$
    for some contact $0<\rho<1$. This in particular implies that for any $\epsilon>0$,
    $\sum_{k=1}^{\infty}{\mathbb{P}}(\mathcal{L}(\bm{\theta}^{(k)})>\epsilon)\leq\sum_{k=1}^{\infty}\frac{{\mathbb{E}}\mathcal{L}(\bm{\theta}^{(k)})}{\epsilon}\leq\frac{\mathcal{L}(\bm{\theta}^{(0)})}{\epsilon}\sum_{k=1}^{\infty}(1-\rho)^{k}=\frac{\mathcal{L}(\bm{\theta}^{(0)})}{\epsilon\rho}<\infty.$
    Now, the Borel–Cantelli lemma gives $\mathcal{L}(\bm{\theta}^{(k)})\to 0$, almost surely.
Appendix B Related Works: Further Details and Discussions
NUQLS shares notable similarities with, and exhibits distinct differences from, several prior works, which are discussed in-depth below.
B.1 Linearized Laplace Approximation (LLA) Framework.
The popular LLA framework (Khan et al., 2019; Foong et al., 2019; Immer et al., 2021; Daxberger et al., 2021) shares a close connection with NUQLS, as both methods fundamentally rely on linearizing the network. However, a subtle yet significant distinction lies in their constructions: LLA begins by obtaining a proper distribution over the parameters and then draws parameter samples from it, while NUQLS bypasses this step and directly targets an approximation of the posterior distribution of the neural network.
As a direct consequence of this, in overparameterized settings, where the Hessian (or its Generalized Gauss-Newton approximation) is not positive definite, the LLA framework necessitates imposing an appropriate prior over the parameters to avoid degeneracy. In sharp contrast, NUQLS directly generates samples from the predictive distribution without introducing any artificial prior, thereby avoiding potential biases that such priors might impose on the covariance structure of the outputs and eliminating the need for additional hyperparameters. Consequently, as the LLA framework corresponds to a Bayesian generalized linear model (GLM), the weight-space vs. function-space duality in GLMs implies that its predictive distribution corresponds to a noisy GP with an NTK kernel. On the other hand, NUQLS leads to a noise-free GP. In interpolating regimes, where the model perfectly fits the data, the noise-free setting of NUQLS appears to be more suitable (Hodgkinson et al., 2023).
Another important consequence of this distinction arises in classification tasks. Due to the Laplace approximation, the LLA framework produces an independent GP for each output of the linearized model. In contrast, NUQLS captures the covariance between outputs, offering a more comprehensive representation of the predictive distribution.
Finally, for regression tasks, NUQLS offers additional flexibility by allowing the variance to be scaled post-hoc by a factor $\gamma$. This enables efficient hyperparameter tuning on a validation set without the need for retraining the model or optimizing a marginal likelihood–a level of flexibility not available in the LLA framework.
NUQLS can be seen as an extension of the “sample-then-optimize” framework for posterior sampling of large, finite-width NNs (Matthews et al., 2017). In this context, the work of Antorán et al. (2022), henceforth referred to as Sampling-LLA, enables drawing samples from the posterior distribution of the LLA in a manner analogous to Algorithm˜1. In this approach, a series of regularized least-squares regression problems are constructed, and the collection of their solutions is shown to be distributed according to the LLA posterior. An EM algorithm is then employed for hyperparameter tuning. In addition to the fundamental differences between NUQLS and LLA-inspired methods mentioned earlier, Sampling-LLA has notable distinctions from Algorithm˜1. First, the objective functions in Sampling-LLA have non-trivial minimum values. As a result, the convergence of SGD for such problems necessitates either a diminishing learning rate (Bubeck et al., 2015), which slows down convergence, or the adoption of variance reduction techniques (Roux et al., 2012; Shalev-Shwartz & Zhang, 2013; Johnson & Zhang, 2013), which can introduce additional computational and memory overhead. By contrast, in overparameterized settings and under the assumptions of Theorem˜3.2, the optimization problem in ˜3 allows interpolation. Consequently, SGD can employ a constant step size for convergence, improving optimization efficiency (Garrigos & Gower, 2023). Second, the inherent properties of the LLA framework, which require a positive definite Hessian or its approximation, necessitate regularizing the least-squares term in the subproblem of Sampling-LLA. This results in a strongly convex problem with a unique solution. Consequently, to generate a collection of solutions, Sampling-LLA constructs a random set of such subproblems, each involving fitting the linearized network to random outputs. These random outputs are sampled from a zero-mean Gaussian, with covariance given by the Hessian of the loss function, evaluated on the data. In contrast, the subproblem of NUQLS, i.e., ˜3, involves directly fitting the training data, and the ensemble of solutions is constructed as a result of random initialization of the optimization algorithm. Hence, the uncertainty captured by NUQLS arises naturally from the variance of solutions in the overparameterized regime, without the need for additional regularization or artificially constructed subproblems.
While the Sampling-LLA method enhances the scalability of LLA, the competing method Variational LLA (VaLLA) (Ortega et al., 2023) offers comparable or superior UQ performance while significantly reducing computation time. VaLLA achieves this by computing the LLA predictive distribution using a variational sparse GP with an NTK. Another competing LLA extension is Accelerated LLA (ELLA), which uses a Nyström approximation of the functional LLA covariance matrix (Deng et al., 2022), and seems to attain similar performance to VaLLA, again at a reduced cost compared to Sampling-LLA. We compare the performance of NUQLS to Sampling-LLA, VaLLA and ELLA in Section˜H.3.
Note that work by (Miani et al., 2024b) approximates an LLA by taking the covariance as the projection onto the null space of the GGN matrix, in order to compute a posterior that retains the same performance as the original network on training data. To compute samples, this work uses alternating projections.
B.2 SNGP
We briefly describe the Spectral-Normalized Neural Gaussian Process (SNGP). SNGP lies at the intersection of feature-space methods and Bayesian methods, and similarly to NUQLS combines a GP with a NN. Specifically, SNGP adds a weight normalization step during training, and then replaces the output layer of NN with a GP that takes the feature extractor of the network as an input. While this method is also not post-hoc, and is not strictly a Bayesian method, we still compare SNGP with NUQLS in Figure˜6 and Table˜11, where we observe superior performance of NUQLS.
B.3 Ensemble Framework.
In Lee et al. (2019), infinitely wide neural networks were shown to follow a GP distribution; however, this GP did not correspond to a true predictive distribution. Building on this, He et al. (2020) introduced a random, untrainable function to an infinitely wide neural network, deriving a GP predictive distribution using the infinite-width NTK. An ensemble of these modified NNs was then interpreted as samples from the GP’s predictive posterior. In contrast, our method demonstrates that trained, finite-width, unmodified linearized networks are inherently samples from a GP with an NTK kernel. While their method, Bayesian Deep Ensembles (BDE), shares some conceptual similarities with ours, we omit it from our main experiments for several reasons. Firstly, the posterior analysis of BDE is valid only in the limit of large model sizes. For smaller datasets and models, the infinite-width NTK differs greatly from the empirical NTK (see Fort et al. (2020)). We can see that in this regime, in both Figure˜4 and Table˜8, that NUQLS outperforms BDE. Secondly, the method is computationally more expensive than DE, due to the need to compute the untrainable function ($\delta(.)$) and tune the scaling hyperparameter for classification. This contradicts our goal to provide a UQ method that is computationally more efficient than the state-of-the-art (SOTA) method DE while maintaining competitive performance. Finally, for larger model sizes, we provide comparison of NUQLS to BDE in Figure˜6 and Table˜11, where we observe that NUQLS outperforms BDE.
In Madras et al. (2019), the authors propose a method called “local ensembles”, which perturbs the parameters of a trained network along directions of small loss curvature to create an ensemble of nearly loss-invariant networks. The uncertainty of the original network is then quantified as the standard deviation across predictions in the ensemble. Building on this approach, Miani et al. (2024a), in a method called Sketched Lanczos Uncertainty (SLU), use the GGN approximation of the Hessian of the loss to identify these directions and introduce a sketched Lanczos algorithm (Meurant, 2006) to efficiently compute them. We compare the performance of NUQLS to SLU in Section˜H.2.
While similar, there are key differences between local ensembles and NUQLS. Local ensembles form a subspace of networks that attain similar loss values, allowing for directions with small but non-zero curvature, potentially encompassing more directions than those with exactly zero curvature. In contrast, when a solution to the linear optimization problem exists, NUQLS creates an ensemble of networks that all attain exactly the same loss. Additionally, NUQLS relies on first-order information to construct this ensemble, whereas local ensembles depend on second-order information.
While the zero-curvature directions of the GGN approximation correspond to the Jacobian’s null space, the local ensembles method also includes directions with small but non-zero curvature. This inclusion introduces a notable distinction between the ensembles generated by local ensembles and those formed by NUQLS. Finally, while local ensembles employ low-rank approximations to compute directions efficiently, such approximations may inadequately represent the Hessian and fail to accurately capture the true curvature structure, as highlighted in Xie et al. (2022).
We also give mention to methods that seek to make DE more efficient, such as BatchEnsemble (BE) (Wen et al., 2020) and SnapshotEnsemble (Havasi et al., 2020). We note that these methods are yet to surpass DE as the SOTA in the literature. However,we compare NUQLS against BE in Figure˜6 and Table˜11, where we see that NUQLS shows superior performance, whilst remaining as scalable as BE. Further, BE requires modification to the structure of the network, and hence is not post-hoc.
B.4 Neural Tangent Kernel Methods
The Procedural-Noise-Correcting (PNC) method of Huang et al. (2023) employs the limiting NTK for infinite-width networks to characterize the noise in the optimization process of training a neural network. It uses this to provide a frequentist confidence interval around test predictions of the neural network, and as it arises from the frequentist framework, PNC recovers statistical guaranetees. In comparison, NUQLS does not rely on the limiting regime of the NTK, but rather employs the feature extraction of the empirical NTK to quantify uncertainty using a Bayesian framework. In Section˜H.1 we compare the performance of NUQLS vs. PNC. Unsurprisingly, given their constructions, we see that PNC outperforms NUQLS when the width of a network is extreme, while for finite-width networks that are fully trained, NUQLS outperforms PNC. This experiment illustrates that the methods are actually complementary, as they both excel in different settings.
Appendix C Uncertainty vs. Prediction
Here we discuss the use of common metrics to measure UQ ability in the BDL community. Generally, NLL, ECE and AUCROC (or the entropic version OODAUC) are used to measure the ability of a method to correctly quantify the uncertainty in a neural network. However, we argue that the use of these metrics for UQ is ill-advised, due to either a misalignment between the metric and the goal of UQ, or due to flawed measurements of uncertainty implicit in the metric. Specifically, we argue that variance over the softmax predictions is the uncertainty in the network, in the same way that the variance over probabilities is the uncertainty in a Dirichlet model. In contrast, NLL and AUCROC (OODAUC) are metrics for prediction, while ECE is a metric based on a flawed surrogate for uncertainty, calibration.
Prediction
While we will argue that uncertainty quantification must involve a computed variance term, and that common UQ metrics are either actually prediction metrics, or are based on a prediction of uncertainty, this discussion also begs the question: what about the purported benefits of BDL to prediction ability? As has been argued in Abe et al. (2022), it is often more economical to simply train a larger capacity model, than to use BDL to improve prediction quality. Hence, we restrict our main focus of NUQLS to its UQ ability. However, for completeness, we also evaluate the novel predictor of NUQLS; these results can be found in the Appendix˜G.
Uncertainty
Uncertainty can be divided into two categories: aleatoric uncertainty, or data uncertainty, and epistemic uncertainty, or model uncertainty. Aleatoric uncertainty arises from intrinsic randomness or noise in the data, and can often not be controlled. In contrast, epistemic uncertainty arises from having many potential models that may fit the data, and the uncertainty in not knowing which model is ‘correct’. As this arises from a lack of knowledge of the original data-mapping, uncertainty in this manner can often be reduced through increased training data, using a larger model, training for longer etc. There are numerous benefits to untangling the aleatoric and epistemic uncertainty (Mukhoti et al., 2021; Huseljic et al., 2021; Chan et al., 2024). We can use the framework of a Dirichlet model, the conjugate prior of a Categorical distribution, to understand how to evaluate these uncertainties in the context of $K$-class classification. Assume there is some subspace $\mathcal{M}$ of the parameter space $\bm{\Theta}\in^{p}$, such that our network ${\bf f}(.;\bm{\theta}^{*})$ attains ‘good’ test performance for any $\bm{\theta}^{*}\in\mathcal{M}$. If it is common across parameter values in $\mathcal{M}$ that ${\bf f}({\bf x};\bm{\theta}^{*})$ has high entropy for some $\bf x$}, i.e. ${\bf f}({\bf x};\bm{\theta}^{*})$ outputs a prediction close to the uniform distribution with high probability across $\bm{\theta}^{*}\in\mathcal{M}$, then we are sure that the outcome is unpredictable in $\bf x$}, i.e. that there is high aleatoric uncertainty. This equates to a Dirichlet model with parameters $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{K})$ where $\bm{\alpha}$ is uniform and of large magnitude. In contrast, assume that over all parameter values in $\mathcal{M}$, the softmax outputs of ${\bf f}({\bf x};\bm{\theta}^{*})$ for a given $\bf x$} have very large variance. This means that no model can agree on the ’correct’ data-mapping, and hence there is very large epistemic uncertainty. This is analogous to when the parameters $\bm{\alpha}$ of a Dirichlet model are small in magnitude. Our method, NUQLS, measures the variance over softmax outputs for high-performing parameters $\bm{\theta}^{*}\in\mathcal{M}$. Hence, we are measuring the epistemic uncertainty of the neural network.
NLL
Due to its probabilistic framework, NLL is often used as a metric for UQ ability. For $K$-class classification, up to a constant the NLL is just the cross-entropy loss
$\text{NLL}(\bm{\theta})=-\sum_{i=1}^{n}\sum_{k=1}^{K}y_{i,k}\log\mu_{i,k},$
where $y_{i,k}$ corresponds to the $k$-th element of the one-hot encoding of the $i$-th label, and $\mu_{i,k}=\text{softmax}({\bf f}({\bf x}_{i};\bm{\theta}))_{k}$. This loss is minimised as $\text{softmax}({\bf f}({\bf x}_{i};\bm{\theta}))_{k}\to\mathbbm{1}_{c_{i}}$, where $c_{i}$ is the correct label for input ${\bf x}_{i}$. While a lower NLL is generally considered to show better UQ ability, we can see that this metric is instead a smooth, probabilistic measurement of accuracy, where proximity of the predictor to the identity function is rewarded with lower NLL. This metric does not measure uncertainty quantification ability.
ECE
A common surrogate for uncertainty in neural networks is the predicted softmax probabilities. While neural networks generally output over-confident probabilities, many methods, such as temperature scaling and many Bayesian methods, attempt to create well-calibrated predictors. A well-calibrated output is one where the predicted softmax probabilities accurately convey the forecasted probability of seeing each of the $K$-classes. For example, of all test points where a network has a maximum predicted probability of $0.9$, we would hope that the network is $90\%$ accurate on these test points. Once a predictor is well-calibrated, the softmax probabilities are then seen to quantify uncertainty. However, there are issues with this:
  1. 1.
    Firstly, a point prediction from a network cannot be interpreted as the uncertainty in the network; uncertainty quantities must accompany a network prediction, for example with a variance term. At best, network confidence can be seen to predict the uncertainty.
  2. 2.
    Secondly, if softmax probability does in fact accurately predict the uncertainty, is this uncertainty due to data uncertainty, or model uncertainty? If a model accurately predicts that there is a $90\%$ chance that class $c$ is correct, does that uncertainty arise due to noise in the data, or some misspecification in the labeling process? Or is it due to the fact that the input is OoD, and thus the model is unsure of itself? Hence, we still require a variance term to ascertain what the epistemic uncertainty is.
  3. 3.
    Finally, calibration is only well-defined for ID test points. To see this, take a network trained on MNIST, and then tested on FashionMNIST. Any network prediction is meaningless on this test set, as the model has been trained for integer labels. The prediction that will minimize the calibration error will be the uniform prediction i.e. the prediction is $0.1,\dots,0.1]\in^{10}$ for all points in the OoD test set. However, the accuracy on OoD test points should be 0. Instead the predicted accuracy is $0.1$, which is what arises from randomly picking one class for each OoD test point. Therefore, the predicted probabilites cannot be directly interpreted as forecasted probability of accuracy for OoD points.
In practice, calibration is tested using ECE. This metric bins predictions, and then computes the average confidence in each bin. This is compared with the average predicted accuracy, and an $l_{2}$ norm is taken over the differences. As was noted in Nixon et al. (2019), there are several issues with ECE. For examples, it is reductive in a multi-class setting, where the bottom K1K-1 probabilities do not contribute meaningfully to the error, and it suffers from the inherent sharpness of neural networks, where the majority of maximum probabilities tend to be very close to $1$, and thus this confidence range is highly over-represented. So we struggle to even tell when a method is well-calibrated.
This issue with ECE can be summarised as follows: confidence is a flawed estimate of the total uncertainty; further, it is hard to tell when a method is well-calibrated. It is thus evident that it is much better to directly model the uncertainty in the model, by calculating the variance over the softmax predictions. Once equipped with this model uncertainty, entropy (or network confidence) can be better trusted to inform of the aleatoric uncertainty.
AUCROC
The final metric commonly used for UQ is AUCROC (OODAUC). This metric uses the top softmax prediction (entropy) as a predictor, to attempt to detect wether a point is ID or OoD. While this is a very useful task, it is still a metric of prediction, as only the mean predictor is used to compute this value. As we have argued, UQ requires a variance term. Miani et al. (2024a) in fact use the variance over logits to compute the AUCROC score, and report this score for their method SLU in comparison to other leading UQ methods. In Section˜H.2 we show that in this metric, we outperform SLU.
VMSP
To combat the issues of the previous metrics, we compare the performance of Bayesian methods using the variance of the maximum softmax predictor, or VMSP. We first provide an explicit definition of VMSP: for a Bayesian method, we generally have a mean predictor $\mu:\mathbb{R}^{d}\to\mathbb{R}^{c}$ and a covariance function $\Sigma:\mathbb{R}^{d}\to\mathbb{R}^{c\times c}$, for example the mean and covariance of the linearized ensemble in the case of NUQLS, that output in the probit space. To compute VMSP for a given test point $\mathbf{x}^{\star}$, we first find $\hat{c}=\text{argmax}_{k}~\mu(\mathbf{x}^{*})_{k}$, where $\mu(\mathbf{x}^{\star})_{k}$ denotes the $k$-th output of $\mu(\mathbf{x}^{\star})$. That is, we find the class that the Bayesian method predicts, given $\mathbf{x}^{\star}$. We then define $\text{VMSP}:=\Sigma(\mathbf{x}^{\star})_{\hat{c},\hat{c}}=\sigma^{2}(\mathbf{x}^{\star})_{\hat{c}}$, that is, the variance of this prediction.
We also introduce a pictorial method for evaluating the performance of Bayesian methods, using VMSP as the correct uncertainty measure. For a given dataset, we want a UQ method to provide low uncertainty for correctly predicted test points, and high uncertainty for incorrectly predicted or OOD test points. We then compare these distributions pictorially using a violin plot, and quantitatively using the median and skew values for the respective distributions. With the addition of a poorly-performing baseline model, we are able to easily compare the ability of UQ models to quantify uncertainty.
Appendix D Guarantees Against Mode Collapse
Given that all linearized models are initialized locally around the trained parameters, one may consider whether it is possible that all linear networks will converge to the same parameter solution, or mode of the linear loss. Fortunately, we are able to show that this will not occur, almost surely. To see this, we note that the solution to ˜3 for each linearized model is given by a unique row-space component (given our assumptions), plus a projection of the initialization ${\bf z}_{i}$ onto the null-space of the Jacobian. For mode-collapse, we would require the projection for two i.i.d initializations ${\bf z}_{1},{\bf z}_{2}\sim N(\mathbf{0},\gamma^{2}\mathbf{I})$ to be equal, or $(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{\dagger}{\mathbf{J}}_{\mathcal{X}})({\bf z}_{1}-{\bf z}_{2})=\mathbf{0}$, where ${\bf z}_{1}-{\bf z}_{2}\neq\mathbf{0}$ with probability one. Hence, ${\bf z}_{1}-{\bf z}_{2}={\mathbf{J}}_{\mathcal{X}}^{\dagger}{\mathbf{J}}_{\mathcal{X}}({\bf z}_{1}-{\bf z}_{2})$, and thus ${\bf z}_{1}-{\bf z}_{2}\in\text{Range}({\mathbf{J}}_{\mathcal{X}}^{T})$, i.e. ${\bf z}_{1},{\bf z}_{2}\in\text{Range}({\mathbf{J}}_{\mathcal{X}}^{T})$. However, $\text{Range}({\mathbf{J}}_{\mathcal{X}}^{T})$ is a low-dimensional subspace of $\mathbb{R}^{p}$, and as we are drawing i.i.d. from a $p$-dimensional normal distribution (i.e. not degenerate), the probability of this event occurring is 0. Hence, we do not need to be concerned with all models converging to the same behaviour.
Appendix E Computational Cost
[…]
Appendix F Hyper-parameter Tuning
NUQLS contains several hyper-parameters: the number of linear networks to be trained, the number of epochs and learning rate of training, and the variance of initialisation, $\gamma$. In this section, we discuss strategies to select optimal hyper-parameters.
F.1 Regression
For regression, and with a sufficiently small learning rate, NUQLS samples from the distribution given in ˜3.3. We can see that the variance of the predictions scales linearly with $\gamma^{2}$. Hence, we use the following framework to tune $\gamma$:
  1. 1.
    To obtain $\bm{\theta}_{s}$ in Algorithm˜1, we initialise our parameters with a very small gamma, e.g. $\gamma=0.01$. This enables (stochastic) gradient descent to converge quickly with a small learning rate.
  2. 2.
    We compute $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},\mathcal{X}_{\text{val}})\}_{s=1}^{S}$, where $\mathcal{X}_{\text{val}}$ is the inputs from a validation set. As per ˜3.3, for each point ${\bf x}\in\mathcal{X}_{\text{val}}$, we compute the mean prediction as $\mu(x)=\text{SampleMean}\left(\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x})\}_{s=1}^{S}\right)$ and the variance as $\sigma^{2}_{\gamma}(x)=\text{SampleVariance}\left(\{\gamma\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},\mathcal{X}_{\text{val}})\}_{s=1}^{S}\right)$ (note the scaling by $\gamma$ for only the variance). We then use these values to compute the ECE across the validation dataset $\mathcal{D}_{\text{val}}$.
  3. 3.
    As coverage of a confidence interval scales linearly with the size of the given standard deviation, and our computed standard deviation scales linearly with $\gamma$, we find that the ECE is convex in $\gamma$. Hence, we employ the Ternary search method (see Algorithm˜2) to find the value $\hat{\gamma}$ that minimizes this validation ECE.
  4. 4.
    For a test point $\bf x$$^{\star}$, our mean prediction and variance is then $\mu({\bf x}^{\star})$ and $\sigma^{2}_{\hat{\gamma}}({\bf x}^{\star})$.
As can be seen in Table˜1 and Table˜9, this framework means that for regression our method is incredibly fast and computes well-calibrated variance values.
Algorithm 2 Ternary Search
$f$: function to minimize, $l$: left boundary of search space, $r$: right boundary of search space, $\delta$: tolerance, iter: iterations.
$i=0$
while $|l-r|\geq\delta$ and i<i<iter do
  $l_{1/3}=l+(r-l)/3$
  $r_{1/3}=r-(r-l)/3$
  if $f(l_{1/3})<f(r_{1/3})$ then
   $l=l_{1/3}$
  else
   $r=r_{1/3}$
  end if
end while
return $(l+r)/2$
F.2 Classification
For uncertainty quantification performance, as measured by VMSP in Figure˜3, we find that as long as $\gamma$ is small, training SGD for a small amount of epochs generally gives small training loss, and hence provides good performance. This means that our method can be computed quite quickly in larger data/model settings.
Appendix G Evaluation of Predictive Mean
While the main focus of NUQLS is the variance term, to compute the epistemic uncertainty of a NN, we would also like to demonstrate the predictive ability of our novel predictive mean, for completeness.
G.1 Image Classification - Predictive
Table 2: Image classification predictive performance, using LeNet5 on MNIST and FashionMNIST (FMNIST). Experiment was run 55 times with different random MAP initialisations to get standard deviation on metrics.
Datasets Method NLL $\downarrow$ ACC $\uparrow$ ECE $\downarrow$ OOD-AUC $\uparrow$ AUC-ROC $\uparrow$ Time (s)
MAP $034\pm0.0$020.002 $\mathbf{0.990\pm}$$\mathbf{0.001}$ $008\pm0.0$010.001 $888\pm0.0$080.008 $886\pm0.0$080.008 $257$
NUQLS $035\pm0.0$020.002 $989\pm0.0$010.001 $\mathbf{0.003\pm}$$\mathbf{0.000}$ $\mathbf{0.930\pm}$$\mathbf{0.026}$ $\mathbf{0.928\pm}$$\mathbf{0.026}$ $106$
MNIST DE $034\pm0.0$040.004 $\mathbf{0.991\pm}$$\mathbf{0.000}$ $011\pm0.0$040.004 $\mathbf{0.932\pm}$$\mathbf{0.009}$ $\mathbf{0.928\pm}$$\mathbf{0.009}$ $2845$
MC-Dropout $044\pm0.0$020.002 $989\pm0.0$000.000 $7\pm0.010.01$ $873\pm0.0$320.032 $871\pm0.0$310.031 $533$
SWAG $\mathbf{0.029\pm}$$\mathbf{0.003}$ $\mathbf{0.991\pm}$$\mathbf{0.000}$ $\mathbf{0.004\pm}$$\mathbf{0.002}$ $902\pm0.0$080.008 $900\pm0.0$080.008 $489$
LLA* $034\pm0.0$020.002 $\mathbf{0.990\pm}$$\mathbf{0.001}$ $008\pm0.0$010.001 $888\pm0.0$080.008 $886\pm0.0$080.008 $45$
VaLLA $034\pm0.0$020.002 $\mathbf{0.990\pm}$$\mathbf{0.001}$ $008\pm0.0$010.001 $889\pm0.0$080.008 $886\pm0.0$080.008 $1583$
MAP $298\pm0.0$070.007 $0.891\pm$$0.3$ $\mathbf{0.006\pm}$$\mathbf{0.001}$ $840\pm0.0$220.022 $804\pm0.0$210.021 $158$
NUQLS $302\pm0.0$060.006 $891\pm0.0$020.002 $\mathbf{0.005\pm}$$\mathbf{0.002}$ $\mathbf{0.904\pm}$$\mathbf{0.007}$ $\mathbf{0.870\pm}$$\mathbf{0.006}$ $89$
FMNIST DE $288\pm0.0$020.002 $896\pm0.0$010.001 $013\pm0.0$010.001 $876\pm0.0$030.003 $836\pm0.0$030.003 $1587$
MC-Dropout $306\pm0.0$070.007 $892\pm0.0$030.003 $026\pm0.0$020.002 $856\pm0.0$210.021 $813\pm0.0$190.019 $291$
SWAG $\mathbf{0.283\pm}$$\mathbf{0.005}$ $\mathbf{0.899\pm}$$\mathbf{0.003}$ $018\pm0.0$020.002 $817\pm0.0$230.023 $783\pm0.0$220.022 $264$
LLA* $298\pm0.0$070.007 $891\pm0.0$030.003 $\mathbf{0.006\pm}$$\mathbf{0.001}$ $841\pm0.0$220.022 $805\pm0.0$210.021 $26$
VaLLA $298\pm0.0$070.007 $891\pm0.0$030.003 $007\pm0.0$010.001 $841\pm0.0$220.022 $805\pm0.0$210.021 $1583$
We compare NUQLS against the MAP NN, DE, MC-Dropout, SWAG, LLA* and VaLLA, on the MNIST and FashionMNIST datasets, using the LeNet5 network. We compare test cross-entropy (NLL), test accuracy (ACC), ECE, and the AUC-ROC measurement for maximum softmax probability (AUC-ROC) and entropy (OOD-AUC) as the detector. The last two metrics evaluate a methods ability to detect out-of distribution points. We display the results in Table˜2. We see that NUQLS performs the best in ECE, AUC-ROC and OOD-AUC, and is competitive in the other metrics, while having the second fastest wall-time.
Appendix H Further Experimental Results
H.1 Comparison to PNC using Confidence Intervals
We use this section to compare the difference in performance of NUQLS vs PNC. To test the difference between the two methods, we tested NUQLS on the confidence intervals problem from Huang et al. (2023). In this experiment, an MLP with a single hidden-layer is trained on $n$ datapoints from $U([0,0.2]^{d})$, where the target is $y=\sum_{i=1}^{d}sin(x_{i})+\epsilon$, and $\epsilon$ is a small noise term. The method is then asked to form a confidence interval around the prediction for ${\bf x}=(0.1,0.1,\dots,0.1)$. This setup is repeated several times, i.e. the network and training data are randomly initialized, and the coverage and average width of the confidence intervals are recorded, as well as the mean prediction. For more details, we refer the reader to Huang et al. (2023). When the width of the network is $32\times n$, and the network is only partially trained with $80$ epochs and a learning rate of $0.01$, NUQLS performs poorly compared to PNC, as can be seen in Table˜3. Note that bolded numbers indicate intervals that have reached or exceeded the expected coverage. Further, smaller width intervals are preferred.
Table 3: Evalaution of coverage (CR) and width (IW) of computed confidence intervals from toy problem, for a neural network with extreme width and partial training. A computed coverage exceeding the expected coverage is bolded. The mean prediction (MP) is also provided.
PNC NUQLS
$95\%$CI (CR/IW) $90\%$CI (CR/IW) MP $95\%$CI (CR/IW) $90\%$CI (CR/IW) MP
d=2,n=128d=2,~n=128 $\mathbf{0.98}/0.0437$ $\mathbf{0.95}/0.0323$ $0.1998$ $0.93/0.0357$ $\mathbf{0.92}/0.0299$ $0.2047$
d=4,n=256d=4,~n=256 $\mathbf{0.98}/0.0411$ $\mathbf{0.95}/0.0304$ $0.3991$ $0.92/0.0596$ $0.86/0.0500$ $0.4084$
Table 4: Evalaution of coverage (CR) and width (IW) of computed confidence intervals from toy problem, for a neural network with finite width and full training. A computed coverage exceeding the expected coverage is bolded. The mean prediction (MP) is also provided.
PNC NUQLS
$95\%$CI (CR/IW) $90\%$CI (CR/IW) MP $95\%$CI (CR/IW) $90\%$CI (CR/IW) MP
d=2,n=128d=2,~n=128 $0.8/0.136$ $0.72/0.0105$ $0.2022$ $\mathbf{0.99}/0.0135$ $\mathbf{0.96}/0.0114$ $0.2012$
d=4,n=256d=4,~n=256 $\mathbf{0.96}/0.0437$ $\mathbf{0.90}/0.0336$ $0.4045$ $\mathbf{0.97}/0.0313$ $\mathbf{0.97}/0.0313$ $0.4030$
d=8,n=512d=8,~n=512 $0.88/0.0740$ $0.88/0.0568$ $0.8078$ $\mathbf{1.00}/0.0667$ $\mathbf{0.98}/0.0559$ $0.8050$
d=16,n=128d=16,~n=128 $0.8/0.1443$ $0.8/0.1108$ $1.6265$ $\mathbf{1.00}/0.1350$ $\mathbf{0.98}/0.1133$ $1.6121$
We note that scaling the width of a network by $32\times n$ is rare in practice, and that such a wide network is difficult to train. If we instead form an MLP with width equal to the number of training points, and increase epochs to $100$ and learning rate to $0.5$, so that the network is properly trained, NUQLS far outperforms PNC, as can be seen in Table˜4. We note that for these experiments we tuned the $\gamma$ hyper-parameter for NUQLS on a small validation set. We conclude that NUQLS and PNC are in fact complementary methods. For infinite-width networks near initialization, PNC performs well while NUQLS struggles. Conversely, for finite-width networks trained to a minimum, NUQLS excels while PNC performs poorly. We believe the latter regime is more representative of practical scenarios, where NUQLS offers significantly better performance.
H.2 Comparison to SLU
Here we compare the performance of our method against the competing SLU method. See Appendix˜B for a discussion of the differences between NUQLS and SLU. Due to the extensive experimental details given in (Miani et al., 2024a), we run certain experiments from this work and report the performance of NUQLS against the SLU results found in (Miani et al., 2024a). In Table˜5, we compare NUQLS against SLU on OoD detection using the AUC-ROC metric, on a smaller single-layer MLP and a larger LeNet model. The MLP is trained on the MNIST dataset, while the LeNet model is trained on the FashionMNIST dataset. For MNIST, the OoD datasets are FashionMNIST, KMNIST, and a Rotated MNIST dataset, where for each experiment run, we compute the average AUC-ROC score over a range of rotation angles (15,30,45,60,75,90,105,120,135,150,165,180)(15,30,45,60,75,90,105,120,135,150,165,180). For FashionMNIST, the OoD datasets are MNIST, and the average Rotated FashionMNIST dataset. The AUC-ROC metric was calculated using the variance of the logits, summed over the classes for each test point, as a score. We observe that NUQLS has a better AUC-ROC value over all OoD datasets, often by a significant margin.
Table 5: Comparing performance of NUQLS against SLU method. Metric given is the AUC-ROC, computed using the variance of the logits, summed over the classes for each test point, as a score. AUC-ROC measures ability of a method to differentiate between ID and OoD points. We see that NUQLS out-performs SLU in these experiments.
Model MLP $p=15k$ LeNet $p=40k$
ID Data MNIST vs FashionMNIST
OoD Data FashionMNIST KMNIST Rotation (avg) MNIST Rotation (avg)
SLU $26\pm0.0$20.02 $42\pm0.0$40.04 $59\pm0.0$20.02 $\mathbf{0.94\pm}$$\mathbf{0.01}$ $74\pm0.0$30.03
NUQLS $\mathbf{0.67\pm}$$\mathbf{0.07}$ $\mathbf{0.79\pm}$$\mathbf{0.02}$ $\mathbf{0.74\pm}$$\mathbf{0.01}$ $\mathbf{0.95\pm}$$\mathbf{0.02}$ $\mathbf{0.91\pm}$$\mathbf{0.01}$
H.3 Comparison to Sampling-LLA, VaLLA and ELLA
We now compare NUQLS against Sampling-LLA, VaLLA and ELLA. Due to issues with convergence, performance, memory usage and package compatibility when either running source code or implementing methods from instructions given in (Antorán et al., 2022) and (Ortega et al., 2023), we instead compare NUQLS against the results for Sampling-LLA, VaLLA and ELLA taken verbatim from (Ortega et al., 2023, Figure 3). We train a $2$-layer MLP, with $200$ hidden-units in each layer, on the MNIST dataset, according to the experimental details given in (Ortega et al., 2023). We then compare the accuracy, NLL, ECE, Brier score and the OOD-AUC metric of NUQLS against those reported for Sampling-LLA, VaLLA and ELLA. The results are shown in Table˜6. We display the mean and standard deviation for NUQLS; as is quoted in (Ortega et al., 2023), the standard deviation for the other methods was below $10^{-4}$ in magnitude, and was thus omitted. We see that NUQLS outperforms Sampling-LLA, VaLLA and ELLA in accuracy, NLL, ECE, and Brier score, and is within one-sixth of a standard deviation of the leading OOD-AUC value, attained by Sampling-LLA. While we cannot directly comment on differences in computation time between our method and these LLA extensions, we were able to run the source code for VaLLA for the experiments in Table˜2, where NUQLS was an order-of-magnitude faster than VaLLA in wall-time. In (Ortega et al., 2023, Figure 3 (right)), we also observe that VaLLA is an order-of-magnitude faster than both Sampling-LLA and ELLA. We also compare with (Ortega et al., 2023, Table 1), where a ResNet20 and ResNet34 model is trained on CIFAR-10. The results are displayed in Table˜7. We see that NUQLS is competitive with all other methods in this setting.
Table 6: Comparing performance of NUQLS against Sampling-LLA, VaLLA and ELLA on MNIST, trained used a 2-layer MLP with $200$ hidden units in each layer, and $\tanh$ activation. Original results taken from (Ortega et al., 2023).
Model ACC NLL ECE BRIER OOD-AUC
ELLA $97.6$ $0.076$ $0.008$ $0.036$ $0.905$
Sampled LLA $97.6$ $0.087$ $0.026$ $0.040$ $\mathbf{0.954}$
VaLLA 100 $97.7$ $0.076$ $0.010$ $0.036$ $0.916$
VaLLA 200 $97.7$ $0.075$ $0.010$ $0.035$ $0.921$
NUQLS $\mathbf{98.0\pm}$$\mathbf{0.1}$ $\mathbf{0.065\pm}$$\mathbf{0.003}$ $\mathbf{0.005\pm}$$\mathbf{0.001}$ $\mathbf{0.031\pm}$$\mathbf{0.001}$ $\mathbf{0.953\pm}$$\mathbf{0.006}$
Table 7: Comparing performance of NUQLS against Sampling-LLA, VaLLA and ELLA on CIFAR-10, with ResNet20 and ResNet32. Original results taken from (Ortega et al., 2023). Purple figures correspond to the top result, while blue figures are the second-best result.
ResNet20 ResNet32
Model ACC NLL ECE ACC NLL ECE
ELLA $92.5$ $0.233$ $0.009$ $93.5$ $0.215$ $0.008$
Sampled LLA $92.5$ $0.231$ $0.006$ $93.5$ $0.217$ $0.008$
VaLLA $92.6$ $0.228$ $0.007$ $93.5$ $0.211$ $0.007$
NUQLS $92.5$ $0.228$ $0.006$ $93.4$ $0.215$ $0.007$
H.4 Comparison to BDE
Figure˜4 displays the performance of BDE on the toy regression problem from Figure˜1. We also compare BDE against NUQLS on several UCI regression tasks in Table˜8.
Figure 4: Comparison of BDE, DE and NUQLS on the toy regression problem from Figure˜1. We can see that the uncertainty of the BDE method is quite small.
Table 8: Comparing performance of NUQLS and BDE on UCI regression tasks. We see that NUQLS outperforms BDE on all tasks.
Dataset Method RMSE $\downarrow$ NLL $\downarrow$ ECE $\downarrow$
Energy BDE $416\pm0.0$390.039 $0.125\pm$$0.212$ $008\pm0.0$050.005
NUQLS $\mathbf{0.047\pm}$$\mathbf{0.006}$ $\mathbf{-2.400\pm}$$\mathbf{0.209}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$
Concrete BDE $714\pm0.0$540.054 $1.563\pm$$0.449$ $063\pm0.0$120.012
NUQLS $\mathbf{0.330\pm}$$\mathbf{0.047}$ $\mathbf{-0.316\pm}$$\mathbf{0.501}$ $\mathbf{0.003\pm}$$\mathbf{0.001}$
Kin8nm BDE $851\pm0.0$370.037 $1.383\pm$$0.582$ $042\pm0.0$120.012
NUQLS $\mathbf{0.252\pm}$$\mathbf{0.005}$ $\mathbf{-0.796\pm}$$\mathbf{0.025}$ $\mathbf{0.000\pm}$$\mathbf{0.000}$
H.5 UCI Regression
We present the results for select UCI regression datasets in Table˜1 in the main body; we present results for several more datasets in Table˜9.
Table 9: Comparing performance of NUQLS, DE, LLA and SWAG on UCI regression tasks. NUQLS performs as well as or better than all other methods, while showing a speed up over other methods; this speed up increases with the size of the datasets. Note that for the Song dataset, the LLA method uses a diagonal covariance structure due to memory constraints: this is denoted as LLA-D.
Dataset Method RMSE $\downarrow$ NLL $\downarrow$ ECE $\downarrow$ Time(s)
Concrete NUQLS $\mathbf{0.330\pm}$$\mathbf{0.047}$ $\mathbf{-0.316\pm}$$\mathbf{0.501}$ $\mathbf{0.003\pm}$$\mathbf{0.001}$ $\mathbf{7.339}~(\mathbf{0.185})$
DE $379\pm0.0$190.019 $\mathbf{-0.574\pm}$$\mathbf{0.098}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $29.047$
LLA $\mathbf{0.333\pm}$$\mathbf{0.050}$ $\mathbf{-0.294\pm}$$\mathbf{0.479}$ $\mathbf{0.003\pm}$$\mathbf{0.002}$ 7.451(0.297)~7.451(0.297)
SWAG $\mathbf{0.334\pm}$$\mathbf{0.050}$ $\mathbf{-0.562\pm}$$\mathbf{0.224}$ $009\pm0.0$060.006 43.416(36.262)43.416~(36.262)
Naval NUQLS $\mathbf{0.049\pm}$$\mathbf{0.012}$ $\mathbf{-2.546\pm}$$\mathbf{0.134}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $\mathbf{11.360}~(\mathbf{0.295})$
DE $076\pm0.0$060.006 $1.761\pm$$0.250$ $093\pm0.0$400.040 $96.570$
LLA $070\pm0.0$220.022 $25.292\pm$$17.570$ $192\pm0.0$290.029 140.724(129.659)140.724~(129.659)
SWAG $1.130\pm$$1.500$ $0.303\pm$$1.091$ $084\pm0.0$220.022 103.727(92.662)103.727~(92.662)
CCPP NUQLS $244\pm0.0$080.008 $885\pm0.0$200.020 $\mathbf{0.000\pm}$$\mathbf{0.000}$ $\mathbf{6.698}~(\mathbf{0.174})$
DE $\mathbf{0.227\pm}$$\mathbf{0.006}$ $\mathbf{-1.009\pm}$$\mathbf{0.041}$ $\mathbf{0.002\pm}$$\mathbf{0.003}$ $79.791$
LLA $243\pm0.0$070.007 $29.420\pm$$4.565$ $163\pm0.0$080.008 38.572(32.048)38.572~(32.048)
SWAG $252\pm0.0$120.012 $849\pm0.0$380.038 $\mathbf{0.001\pm}$$\mathbf{0.002}$ 73.357(66.833)~73.357~(66.833)
Wine NUQLS $\mathbf{0.789\pm}$$\mathbf{0.042}$ $\mathbf{0.284\pm}$$\mathbf{0.066}$ $\mathbf{0.001\pm}$$\mathbf{0.000}$ $\mathbf{1.164}~(\mathbf{0.115})$
DE $\mathbf{0.789\pm}$$\mathbf{0.041}$ $\mathbf{0.320\pm}$$\mathbf{0.109}$ $\mathbf{0.001\pm}$$\mathbf{0.001}$ $13.241$
LLA $\mathbf{0.792\pm}$$\mathbf{0.041}$ $1.012\pm$$0.182$ $009\pm0.0$040.004 1.389(0.340)1.389~(0.340)
SWAG $\mathbf{0.798\pm}$$\mathbf{0.038}$ $\mathbf{0.367\pm}$$\mathbf{0.103}$ $005\pm0.0$030.003 12.856(11.807)12.856~(11.807)
Yacht NUQLS $\mathbf{0.042\pm}$$\mathbf{0.013}$ $\mathbf{-1.561\pm}$$\mathbf{2.319}$ $\mathbf{0.012\pm}$$\mathbf{0.010}$ $\mathbf{3.390}~(\mathbf{0.164})$
DE $0.647\pm$$0.121$ $2.032\pm$$0.349$ $\mathbf{0.016\pm}$$\mathbf{0.008}$ $40.132$
LLA $\mathbf{0.043\pm}$$\mathbf{0.014}$ $\mathbf{-2.733\pm}$$\mathbf{0.468}$ $\mathbf{0.011\pm}$$\mathbf{0.006}$ 3.403(0.177)3.403~(0.177)
SWAG $\mathbf{0.044\pm}$$\mathbf{0.014}$ $2.565\pm$$0.118$ $067\pm0.0$250.025 19.408(15.822)19.408~(15.822)
Song NUQLS $\mathbf{0.839\pm}$$\mathbf{0.014}$ $646\pm0.0$560.056 $1\pm0.0000.0$$0$ $\mathbf{295.058}~(\mathbf{91.673})$
DE $\mathbf{0.846\pm}$$\mathbf{0.006}$ $\mathbf{0.180\pm}$$\mathbf{0.013}$ $5\pm0.0000.0$$0$ $2562.789$
LLA-D $\mathbf{0.851\pm}$$\mathbf{0.029}$ $456\pm0.0$930.093 $\mathbf{0.000\pm}$$\mathbf{0.000}$ 413.814(210.429)413.814~(210.429)
SWAG $845\pm0.0$020.002 $680\pm0.0$620.062 $003\pm0.0$010.001 3477.825(3274.440)3477.825~(3274.440)
H.6 eNUQLS
In Figure˜5 we demonstrate the performance of an ensembled version of NUQLS, eNUQLS. This method is similar to a Mixture of Laplace Approximations (Eschenhagen et al., 2021). We observe excellent separation of the variances between correct and incorrect/OoD test groups for eNUQLS, especially for CIFAR-10 on ResNet9. Note that there is significant cost to ensembling our method, and we provide this figure simply to illustrate performance capacity.
Figure 5: Violin plot of VMSP, with an ensembled version of NUQLS, eNUQLS, included.
H.7 Additional VMSP Results
Figure 6: Violin plot of VMSP, for (top left) ResNet50 SVHN, (top right) ResNet50 ImageNet, (bottom left) WRN-34-1 FMNIST, and (bottom right) ResNet9 FMNIST.
We use this section to expand the empirical evaluation of NUQLS. In the top left of Figure˜6, we observe that on a ResNet50 model trained on SVHN, with CIFAR-10 as the OoD test data, NUQLS outperforms all other methods. In the top right of Figure˜6, we evaluated NUQLS on the ImageNet dataset in order to display the scalability of our method to larger datasets. We employed the pre-trained weights for a ResNet50, as found on torch.hub, and used the ImageNet-o dataset as our OOD test set (Hendrycks et al., 2021). Due to resource budget constraints, we were only able to compare NUQLS against a baseline method, though future research could include a comparison of other methods against NUQLS, using VMSP, on ImageNet. As can be seen, we see excellent separation between correct and incorrect predictions, while only adequate separation between correct and OOD points. We note from Bitterwolf et al. (2023) that approximately $21\%$ of the images in ImageNet-o are actually in-distribution; hence, we see that NUQLS correctly quantifies the variance of these test sets. We also provide comparison of NUQLS with additional competing methods. In the bottom left of Figure˜6, we compare NUQLS against the BE method. Note that BE requires modification to the structure of a neural network; hence, we compare BE on a modified WideResNet-34-1, of which an implementation of the correct modifications existed (Franchi et al., 2023), trained on FashionMNIST (with MNIST as OOD). We again observe that NUQLS shows greater separation than BE for VMSP across correct predictions versus incorrect or OOD predictions. Finally, in the bottom right of Figure˜6, we compare NUQLS against further competing methods, specifically BDE, SGLD (Welling & Teh, 2011) (which we include to demonstrate the inferior performance of MCMC methods in this regime) and SNGP on ResNet9 trained on FMNIST (with MNIST as OOD); note SNGP requires modifications to the training procedure. We see that BDE performs similarly to DE, and SLGD performs similarly to SWAG, which is unsurprising considering their respective constructions. We also observe that NUQLS still performs the strongest across all methods using the VMSP violin plots. The respective median and skew values for Figure˜6 are displayed in Table˜10 and Table˜11; these values also evidence the strong performance of NUQLS.
H.8 Image Classification Skew and Median
The median and sample skew for the VMSP in Figure˜3 and Figure˜6 is found in Table˜10 and Table˜11.
Table 10: Sample median and sample skew of variance from Figure˜3 and Figure˜6. IDC = ID correct, IDIC = ID incorrect, FMNIST = FashionMNIST. The median and skew for a method is compared against a baseline method. We expect positive skew for the ID correct (IDC) variances, and a negative skew for the ID incorrect (IDIC) and OoD variances. If the median or skew for a method is worse than the corresponding baseline, then it is written in gray.
Median NUQLS DE SWAG MC LLA BASE
ResNet9 IDC $\downarrow$ $4.76\times 10^{-15}$ $1.31\times 10^{-5}$ $4.64\times 10^{-7}$ $5.05\times 10^{-7}$ $9.02\times 10^{-10}$ $0.020$
[…]
ResNet50 IDC $\downarrow$ $0.00$ $8.29\times 10^{-8}$ $0.018$ $1.49\times 10^{-9}$ $1.70\times 10^{-7}$ $0.019$
[…]
CIFAR-100 IDIC $\uparrow$ $0.211$ $0.0624$ $0.101$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{5.93\times 10^{-3}}$ $0.0225$ $0.020$
[…]
ResNet50 IDC $\downarrow$ $8.61\times 10^{-15}$ $3.59\times 10^{-6}$ $1.71\times 10^{-6}$ $2.18\times 10^{-8}$ $1.61\times 10^{-6}$ $0.020$
SVHN IDIC $\uparrow$ $0.217$ $0.0569$ $0.0302$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{3.96\times 10^{-4}}$ $0.0241$ $0.020$
OoD $\uparrow$ $0.233$ $0.127$ $0.0789$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{7.51\times 10^{-4}}$ $0.0252$ $0.020$
ResNet50 IDC $\downarrow$ $0.00$ - - - - $0.020$
ImageNet IDIC $\uparrow$ $0.232$ - - - - $0.020$
OoD $\uparrow$ $0.100$ - - - - $0.020$
Sample Skew
ResNet9 IDC $\uparrow$ $2.4$ $3.51$ $3.72$ $4.92$ $6.46$ $1.01$
[…]
ResNet50 IDC $\uparrow$ $2.97$ $2.68$ $4.02$ $4.55$ $5.14$ $1.14$
[…]
ResNet50 IDC $\uparrow$ $2.2$ $2.42$ $3.52$ $6.91$ $4.15$ $0.989$
[…]
ResNet50 IDC $\uparrow$ $1.52$ - - - - $1.21$
ImageNet IDIC $\downarrow$ 0.88-0.88 - - - - $1.10$
OoD $\downarrow$ $0.179$ - - - - $0.985$
Table 11: Sample median and sample skew of variance from Figure˜6. IDC = ID correct, IDIC = ID incorrect, FMNIST = FashionMNIST. The median and skew for a method is compared against a baseline method. We expect positive skew for the ID correct (IDC) variances, and a negative skew for the ID incorrect (IDIC) and OoD variances. If the median or skew for a method is worse than the corresponding baseline, then it is written in gray.
Median NUQLS BE BDE SGLD SNGP BASE
WRN-34-1 IDC $\downarrow$ $0.00$ $8.31\times 10^{-5}$ - - - $0.020$
FMNIST IDIC $\uparrow$ $0.178$ $0.045$ - - - $0.020$
OoD $\uparrow$ $0.173$ $0.107$ - - - $0.020$
ResNet9 IDC $\downarrow$ $4.76\times 10^{-15}$ - $3.35\times 10^{-20}$ $6.79\times 10^{-8}$ $9.19\times 10^{-7}$ $0.020$
[…]
Sample Skew
WRN-34-1 IDC $\uparrow$ $3.1$ $2.8$ - - - $1.01$
FMNIST IDIC $\downarrow$ 0.342-0.342 $0.944$ - - - $1.11$
OoD $\downarrow$ 0.0875-0.0875 0.272-0.272 - - - $1.11$
ResNet9 IDC $\uparrow$ $2.4$ - $6.46$ $5.49$ $29.8$ $0.93$
[…]
Appendix I Experiment Details
All experiments were run either on an Intel i7-12700 CPU (toy regression), or on an H100 80GB GPU (UCI regression and image classification). Where multiple experiments were run, mean and standard deviation were presented.
I.1 NTK Convergence
For this experiment, 2 randomly sampled sets of $100$ Gaussian inputs of dimension $5$, and $100$ scalar Gaussian targets, were used as training points and test points respectively. The weights of the MLP were initialised to $\mathcal{N}(0,1)$; no bias terms were used. To train, GD was employed with a Nesterov momentum parameter of $0.9$, a learning rate of $0.1$, and $5000$ epochs. For NUQLS, a learning rate of $0.1$ was used, and $\gamma$ was set to $1$. Error bars on Figure˜2 are the $95\%$ sample confidence interval.
I.2 Toy Regression
We use a 1-layer MLP, with a width of $50$ and SiLU activation. For the maximum a posteriori (MAP) network, we train for $10000$ epochs, with a learning rate of $0.001$, using the Adam optimizer and the PyTorch polynomial learning rate scheduler, with parameters total_\_iters == epochs, power =0.5=~0.5. For DE, each network in the ensemble outputs a heteroskedastic variance, and is trained using a Gaussian NLL, with $2000$ epochs and a learning rate of $0.05$. We combine the predictions of the ensembles as per (Lakshminarayanan et al., 2017). Both DE and NUQLS use $10$ realizations. The $\gamma$ hyper-parameter in NUQLS is set to $5$, and each linear realization is trained for $1000$ epochs with a learning rate of $0.001$, using SGD with a momentum parameter of $0.9$. In SWAG, the MAP network is trained for a further $10000$ epochs, using the same learning rate, and the covariance is formed with a rank-$10$ approximation. A prior precision of $0.1$ and $1$ is used for LLA and LA respectively, as well as the full covariance matrix. The variational inference method used is Bayes By Backprop (Blundell et al., 2015), as deployed in the Bayesian Torch package (Krishnan et al., 2022). The prior parameters are ($\mu=0,\sigma=1$), and the posterior is initialized at ($\mu=0,\rho=-3$). For SWAG, LLA, LA and VI, $1000$ MC sample were taken at test time. These design choices gave the best performance for this problem.
I.3 UCI Regression
Table 12: Training procedure for UCI regression results in Table˜1 and Table˜9.
NN Energy Concrete Kin8nm Naval CCPP Wine Yacht Protein Song
Learning Rate $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$
Epochs $1500$ $1000$ $500$ $150$ $100$ $100$ $1000$ $250$ $50$
Weight Decay $10^{-5}$ $10^{-5}$ $10^{-5}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ $10^{-5}$ $10^{-4}$ 0
Optimizer Adam Adam SGD SGD Adam SGD Adam SGD SGD
Scheduler PolyLR PolyLR None None PolyLR None PolyLR None None
MLP Size $[150]$ $[150]$ $[100,100]$ $[150,150]$ $[100,100]$ $[100]$ $[100]$ $[150,200,150]$ $[1000,1000$
,500,50],500,50]
NUQLS
Learning Rate $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-3}$
Epochs $150$ $100$ $50$ $15$ $10$ $10$ $100$ $25$ $10$
DE
Learning Rate $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$
Epochs $1500$ $300$ $100$ $100$ $100$ $100$ $1000$ $250$ $50$
Optimizer Adam Adam Adam Adam Adam Adam Adam Adam Adam
Scheduler None None Cosine None None None Cosine None None
Experiment
No. experiments $10$ $10$ $10$ $10$ $10$ $10$ $10$ $10$ 33
We now provide the experimental details for the UCI regression experiments (as seen in Table˜1 and Table˜9). For each dataset, we ran a number of experiments to get a mean and standard deviation for performance metrics. In each experiment, we took a random $70\%/15\%/15\%$ split of the dataset for training, testing, and validation. The training hyper-parameters for the MAP, DE and NUQLS networks, size of the MLP used, and the number of experiments conducted for each dataset can be found in Table˜12. Reported time for NUQLS, LLA and SWAG includes the training time for the original NN, with the time to run the method given in brackets.
[…]
  • NUQLS: The linear networks were trained using (S)GD with Nesterov momentum parameter $0.9$. For all datasets except for Song, the full training Jacobian could be stored in memory; this made training extremely fast. For the Song dataset, we trained all linear networks in parallel, by explicitly computing the gradient using JVPs and VJPs. The number of linear networks used was $10$ across all datasets, and the $\gamma$ hyper-parameter was kept at $0.01$.
  • DE: Each member of the ensemble output a separate heteroskedastic variance, and was trained to minimise the Guassian negative log likelihood. The ensemble weights were also initialized as Xavier normal, and bias as standard normal. The number of ensemble members was kept at $10$.
  • LLA: LLA requires two parameters for regression: a dataset noise parameter, and a prior variance on the parameters (Foong et al., 2019). To find the noise parameter, a grid search over $10$ values on a log-scale between 1e21e-2 and $1e2$ was used to find the noise that minimized the Gaussian likelihood of the validation set, with the LLA mean predictor as the Gaussian mean. The same grid search was used to find the prior variance, in order to minimize the expected calibration error (ECE) of LLA on the validation set. For the Protein dataset, a Kronecker-Factored Curvature (KFAC) covariance structure was used (Immer et al., 2021), and for the Song dataset a diagonal covariance structure was used. For all other datasets, LLA used the full covariance structure. The predictive distribution was computed with $1000$ MC samples.
  • SWAG: SGD was used, and the learning rate and number of epochs was kept the same as the NN. We used a grid-search for weight decay, over the values $[0,0.005,0.00005]$, to minimize the ECE on the validation set. The rank of the covariance matrix was $10$ for all datasets, and the predictive distribution used $1000$ MC samples.
I.4 Image Classification
We display the training procedure in Table˜13 for both Figure˜3 and Table˜2. For MNIST and FashionMNIST, we took a $5:1$ training/validation split of the training data. For CIFAR-10, we simply used the test data as a validation set. For CIFAR-10, random horizontal crop and flip on the training images was used as regularization.
  • NN: We chose the training procedures to provide the best MAP performance. All networks have weights initialized as Xavier uniform. For SGD, a momentum parameter of $0.9$ was used. For the Cosine Annealing learning rate scheduler, the maximum epochs was set to the training epochs.
  • NUQLS: The number of samples was kept to $10$ for all datasets.
  • DE: Similary, $10$ ensemble members were used for all datasets.
  • MC-Dropout: Dropout was applied to the network before the last fully connected layer. The dropout probability was set to $0.1$ (a larger probability of $0.25$ was also used, but it did not change the result). At test time, $100$ MC-samples were taken.
  • SWAG: The network was trained for a further $1\times$ training epochs. For CIFAR-$100$ and SVHN, learning rate was kept the same as the original NN; for other datasets, a larger learning rate of $1e2\timesNN$ learning rate was used. The covariance rank was set at $10$. At test time, $100$ MC-samples were taken.
  • LLA*: We used a last-layer KFAC approximation to the covariance. The prior precision was found through a grid search over $20$ values on a log scale from 1e21e-2 to $1e2$, using the probit approximation to the predictive, and a validation set. This configuration for LLA is what is recommended in Daxberger et al. (2021). We used $1000$ samples, to remedy the large amount of approximations used.
  • VaLLA: We used the implimentation found in (Ortega et al., 2023). We kept nearly all hyper-parameters the same as in the MNIST and FashionMNIST experiments in (Ortega et al., 2023); however, we reduced the number of iterations to $5000$, due to time constraints. We were unable to run this implimentation for ResNet9 or ResNet50, due to an out-of-memory error.
Table 13: Training procedure for image classification results in Table˜2, Figure˜3, and Figure˜6.
MAP LeNet5 MNIST Lenet5 FMNIST ResNet9 FMNIST ResNet50 CIFAR10 ResNet50 CIFAR100 ResNet50 SVHN ResNet50 ImageNet
Learning Rate $5\times 10^{-3}$ $5\times 10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ $10^{-2}$ -
Epochs $35$ $35$ $10$ $200$ $200$ $50$ -
Weight Decay $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $5\times 10^{-4}$ $5\times 10^{-4}$ -
Batch Size $152$ $152$ $100$ $128$ $128$ $128$ -
Optimizer Adam Adam Adam SGD SGD SGD -
Scheduler Cosine Cosine Cosine Cosine Cosine Cosine -
Accuracy $99\%$ $90\%$ $\% 92.592.5\%$ $75\%$ $87\%$ $76\%$
NUQLS
Learning Rate $10^{-2}$ $10^{-2}$ $10^{-1}$ $10^{-1}$ $10^{-2}$ $10^{-2}$ $10^{-2}$
Epochs $10$ $10$ 22 $50$ $10$ 22 22
Batch Size $152$ $152$ $50$ $128$ $128$ $256$ $56$
$\gamma$ 11 $0.7$ $0.1$ $0.01$ $0.05$ $0.05$ $0.7$
I.5 Confidence Intervals
For the confidence intervals experiments in Table˜4 and Table˜3, we used $100$ repeats for $d\leq 4$ and $50$ repeats for $d\geq 8$. We used a multiplier of $1.5$ on all NUQLS variance values, as we note that the search strategy for $\gamma$ for regression only employs ECE; this does not reward confidence intervals that exceed the expected confidence, and hence computed confidence intervals may be conservative in width.
I.6 Comparison Experiments
I.6.1 BDE
For Figure˜4 and Table˜8, BDE was implimented following details given in (He et al., 2020). To make comparison with NUQLS fair, each ensemble member of BDE did not output a separate hetero-skedastic variance term. Instead, the variance over predictions was taken as the variance. Further, the large computational cost of BDE relative to NUQLS meant that it is unrealistic for comparison to allow BDE to tune hyper-parameters through a grid-search, as this would take the run time of BDE (which is slightly greater than that of DE) severely out-of-proportion to NUQLS. For Figure˜4, the noise-parameter was given as the noise in the data, $9$. For Table˜8, standard Gaussian noise was taken a priori. The ensemble members were trained with the same training scheme as the original NN in both Figure˜4 and Table˜8. For both experiments, $10$ ensemble members were used. For classification, we again followed the implementation details given in He et al. (2020) to the best of our ability. We employed the same hyperparameters as DE/MAP, and used $10$ ensemble members.
I.6.2 SLU
For the results in Table˜5, we copied the training details for (Miani et al., 2024a, Table 3.); we point the reader to (Miani et al., 2024a, Appendix D.1) for the exact details. For NUQLS, we used the following hyper-parameters for both MNIST and FashionMNIST: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ 11.
I.6.3 Sampling-LLA, VaLLA and ELLA
For the results in Table˜6, we used the training details for (Antorán et al., 2022, Figure 3. (left)); again, we point the reader to the details given in (Antorán et al., 2022, Appendix F.1). For NUQLS, we used the following hyper-parameters for MNIST: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ $0.25$. For ResNet20 the following hyper-parameters were used: Epochs $1$, samples $100$, learning rate $0.0001$, batch size $152$, $\bm{\gamma}$ $0.01$. For ResNet32 the following hyper-parameters were used: Epochs $1$, samples $150$, learning rate $0.0001$, batch size $152$, $\bm{\gamma}$ $0.01$.
I.6.4 BE
We employed the WideResNet architecture from Franchi et al. (2023). We used a depth of $10$ and a widening factor of $1$. The MAP network was trained with: Epochs $20$, learning rate $0.005$, weight decay $0.0001$, batch size $152$, the Adam optimizer and the Cosine learning rate scheduler. BE employed the same hyperparameters, yet with $1.5\timesEpochs$, to aid with training. BE used $5$ ensemble members, in order to match the computational budget of NUQLS. The NUQLS hyperparameters were: Epochs $10$, samples $10$, learning rate $0.01$, batch size $152$, $\bm{\gamma}$ 11.
I.6.5 SNGP
We employed the Lightning UQ Box (Lehmann et al., 2025) implementation of SNGP for our experiments. The same hyperparameters as the MAP network were used. Note that we also tested SNGP on ResNet50 SVHN, however SNGP was unable to train successfully, even after trying $3$ learning rates and $2$ optimizers. Hence, we did not include the results for ResNet50 SVHN.
I.6.6 SGLD
We have implemented the SGLD method from Welling & Teh (2011) using the Lightning UQ Box package (Lehmann et al., 2025) as the basis of the code. We then amended this code to include the learning rate scheduler from Welling & Teh (2011). We followed Maddox et al. (2019), and initialized the SGLD trajectory from the weights of the trained network. We also copied the learning rate from Maddox et al. (2019), and used the same weight-decay as the original network for the SGLD prior. However, in contrast to Maddox et al. (2019), we took the noise-factor scaling to be $10^{-3}$ instead of $5\times 10^{-4}$, as we found that this gave better performance. We sampled $100$ epochs from the posterior, using a batch size of $100$.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We claim that our method approximates the posterior of an NTK-GP, and provide theoretical and empirical proof in Section˜3 and Section˜4.1 respectively. We also claim that we outperform competing methods in performance and efficiency, which can be seen in Section˜4 and Appendix˜H.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Discussion of our limitations can be found in the "Limitations" subsection of the Conclusion, Section˜5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We contain two provable results, Lemma˜3.1 and Theorem˜3.2. The proofs can be found in Appendix˜A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All code for the experiments has been included with the submission. All experimental details have been described in Appendix˜I.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. 1.
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. 2.
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. 3.
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. 4.
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Link to code, as well as a downloadable package of our method, can be found at the top of Section˜4.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All training and test details are found in Appendix˜I.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: UCI Regression experiments (Tables˜1, 9 and 8 and the empirical convergence figure (Figure˜2) have error bars included, and are defined in Figure˜2 and in Section˜I.3. Violin plots do not have error bars, as they are generally not reported with error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Experimental resources are listed at the top of Appendix˜I.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: After reading through the NeurIPS Code of Ethics, we believe that we have conformed to this code.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the benefits of improved UQ methods in our Introduction and Abstract. We are not aware of any direct negative impacts of our work (though secondary negative impacts may exist).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We did not use any high risk models or datasets, and instead only use highly common models and datasets that are commonly used in the UQ literature.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Any prior code for related methods that has been used has been referenced in the linked GitHub repository for our code. Any related methods for comparison has been properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our method has been released as package, with examples given for how to correctly use our method.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The development of our method did not involve the use of LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.